### v2

# Beyond the noise: intrinsic dimension estimation with optimal neighbourhood identification

Abstract
The Intrinsic Dimension (ID) is a key concept in unsupervised learning and feature selection, as it is a lower bound to the number of variables which are necessary to describe a system. However, in almost any real-world dataset the ID depends on the scale at which the data are analysed. Quite typically at a small scale, the ID is very large, as the data are affected by measurement errors. At large scale, the ID can also be erroneously large, due to the curvature and the topology of the manifold containing the data. In this work, we introduce an automatic protocol to select the sweet spot, namely the correct range of scales in which the ID is meaningful and useful. This protocol is based on imposing that for distances smaller than the correct scale the density of the data is constant. In the presented framework, to estimate the density it is necessary to know the ID, therefore, this condition is imposed self-consistently. We derive theoretical guarantees and illustrate the usefulness and robustness of this procedure by benchmarks on artificial and real-world datasets.
footnotetext: *antonio.dinoia(at)stat.math.ethz.ch. 1Seminar for Statistics, Department of Mathematics, ETH Zurich. 2Faculty of Economics, Euler Institute, Università della Svizzera italiana. 3International School for Advanced Studies (SISSA). 4Department of Translation and Language Sciences, Universitat Pompeu Fabra. 5Banca d’Italia. The views and opinions expressed in this paper are those of the authors and do not necessarily reflect the official policy or position of Banca d’Italia. 6Department of Science and High Technology, University of Insubria.
Keywords: high-dimensional data, Poisson process, dimensionality reduction, intrinsic dimension, nearest-neighbour methods.
1 Introduction
The study of high-dimensional data is a fundamental task in modern statistics, as many real-world datasets are often characterized by a large number of variables, a fact that may hinder their interpretation and analysis. In most high-dimensional applied settings, a subset of the variables or non-linear combinations of them are often sufficient to describe the dataset. The Intrinsic Dimension (ID) represents the minimum number of independent variables that allow to locally reproduce the data with minimal information loss. This quantity is an integral part of several data science methods, from dimensionality reduction, where it can provide a meaningful target (Rozza et al.,, 2012; Gliozzo et al.,, 2024; Van Der Maaten et al.,, 2009), to density estimation or clustering, where it can be used to improve the accuracy and robustness of existing algorithms (Rodriguez et al.,, 2018). Furthermore, the ID has become increasingly used as a stand-alone data-driven observable to quantify the complexity of datasets. Usage of ID estimation in this sense can be found in different scientific disciplines such as in biology (Facco et al.,, 2019), genomics (Macocco et al.,, 2023), quantum physics (Mendes-Santos et al., 2021a, ; Mendes-Santos et al., 2021b, ), computer science (Ansuini et al.,, 2019; Valeriani et al.,, 2024), image analysis (Pope et al.,, 2021). A further context where the intrinsic dimension is recently arising is in the theoretical analysis of various statistical methods, in particular regarding convergence rates when data are assumed to have a low-dimensional structure, see e.g. Kpotufe, (2011), Nakada and Imaizumi, (2020), Rosa et al., (2024), Berenfeld et al., (2022), Rosa and Rousseau, (2024), Tang et al., (2024).
Many different, but practically equivalent, ID definitions have appeared in the literature. For example, it has been defined as the dimension of the support of the density which is generating the data. A comprehensive overview of these definitions can be found in (Lee et al.,, 2015). To estimate the ID, many approaches have been proposed, including fractal (e.g. Erba et al.,, 2019), topological (e.g. Block et al.,, 2022), nearest-neighbour (NN), and graph-based methods (e.g. Serra et al.,, 2017). We build on some very recently proposed NN ID estimators (Facco et al.,, 2017; Levina and Bickel,, 2004; Macocco et al.,, 2023; Denti et al.,, 2022), as they are statistically well-grounded and hence allow for uncertainty quantification and exact or asymptotic distributional results.
Despite the favourable statistical foundations, all NN-based ID estimators possess a set of arbitrary free parameters that strongly control the scale at which one observes the data or, equivalently, the level of the dataset decimation (Facco et al.,, 2017; Denti et al.,, 2022). This often leads to unclear choices for the right scale or the proper decimation level to consider, which can easily produce significantly biased estimates. To address this issue, we propose a robust adaptive estimation methodology with provable theoretical guarantees. The methodology is formulated in a likelihood-based framework and further extended to its Bayesian counterpart. As we will show, our proposal brings three main advantages. First of all, it allows for controlling that the locally constant density assumption is satisfied, which is a crucial step when employing any NN-based methodology both in ID estimation as well as in density estimation. The second advantage is related to the fact that many real datasets are affected by very noisy measurements. We will show that our proposal allows us to overcome this issue by automatically escaping the noisy neighbourhoods of the observations. Finally, and most importantly, our methodology allows for the automatic selection of the scale at which the researcher should look at the data. This is done in a point-wise adaptive fashion, which avoids the need for an arbitrary choice of the scale which is, conversely, typically kept constant across the entire dataset. The optimal scale identification is a fundamental and often overlooked feature/passage. Real-world datasets can often be described by hierarchical structures characterized by a density varying from region to region. For such reasons, providing an ID estimate without referring to its scale can be misleading. Furthermore, this scale selection turns out to be crucial to choosing the correct noise-to-signal ratio, which would otherwise be completely unknown. In addition, our estimation framework is computationally feasible and allows for a straightforward uncertainty quantification thanks to the asymptotic normality of the estimator.
The remainder of the paper is organized as follows. In Section 2, we introduce our robust adaptive ID estimator. In Section 3, we illustrate our methodology on some simulated and real datasets, comparing its performance with non-adaptive procedures. Some concluding remarks are offered in Section 4. The Appendix reports theoretical guarantees for the proposed methodology, additional experiments and some extensions of the method.
2 Methodology
2.1 Binomial Intrinsic Dimension Estimator (BIDE)
Here we introduce our modelling framework which builds on Macocco et al., (2023) where the authors propose the Intrinsic Dimension estimator for Discrete Datasets (I3D) specifically tailored for estimating the ID in discrete spaces. Let $(\Omega,\mathcal{A},P)$ be a probability space and let $(\mathcal{M}^{D},\Sigma,\mu)$ be a $D$-dimensional measure space. Let us consider $X_{1},\dots X_{n}$ independent copies of a random vector $X:\Omega\to\mathcal{M}^{D}$ sampled nearby a $d$-dimensional smooth manifold embedded in $\mathcal{M}^{D}$. Consider a measurable subset $A\subset\mathcal{M}^{D}$ with measure $\mu(A)$. Let us consider a Poisson point process with intensity function $h:\mathcal{M}^{D}\to\mathbb{R}_{+}$ that is constant and equal to $\rho$ on the subset $A$, thus, the number of data points in $A$ denoted by $k_{A}$ is a Poisson random variable with parameter $\rho\mu(A)$ and probability mass function (pmf) $p_{k_{A}}$ given by
$p_{k_{A}}(x)=\frac{(\rho\mu(A))^{x}}{x!}\exp(-\rho\mu(A)).$
Now, let $A\subset B\subset\mathcal{M}^{D}$ where $B$ is further measurable subset and assume $h$ is constant and equal to $\rho$ also on $B$, we get the distribution of the number of points in $B\setminus A$, namely, $k_{B}-k_{A}\sim\mathrm{Poisson}(\rho\mu(B\setminus A))$ and the conditional distribution $k_{A}|k_{B},p\sim\mathrm{Binomial}(k_{B},p)$ with $p=(\rho\mu(A))/(\rho\mu(B))$. Thus, the underlying Poisson process is assumed to be homogeneous only locally, on the scale of the $B$ set. Note that the intensity function can be regarded as the non-normalized measure underlying the data-generating process, therefore the local homogeneity assumption of the Poisson process is often regarded as a constant density assumption, assuming it exists with respect to $\mu$.
At this stage, we can take into account the nature of the space $\mathcal{M}^{D}$. We set $\mathcal{M}^{D}\equiv\mathbb{R}^{D}$ equipped with the Euclidean metric, while $\Sigma$ is the Borel $\sigma$-field and $\mu$ is the Lebesgue measure on $\mathbb{R}^{D}$ which is locally equivalent to the Lebesgue measure on $\mathbb{R}^{d}$ i.e. defined on the tangent space to the low-dimensional manifold embedded in $\mathbb{R}^{D}$. Let $A$ and $B$ open balls centred at the same point on the tangent space of such manifold and s.t. $A\subset B$, then $p=(t_{A}/t_{B})^{d}=\tau^{d}$ where $t_{A}$ and $t_{B}$ are the radii of the two balls. Here, $d$ is the ID, i.e. the dimension of the manifold and we assume that the random vectors $X_{1},\dots,X_{n}$ are sampled nearby such manifold. Let $k_{A,i}$ and $k_{B,i}$ be, respectively, the number of points in the open balls $A$ and $B$ centred at unit $X_{i}=x_{i}$, where we do not count $x_{i}$ since we condition on it. If $k_{A,1},\dots,k_{A,n}$ are independent conditionally on $k_{B,1},\dots,k_{B,n},\tau$, the likelihood of $d$ reads
$L(d|k_{B,1},\dots,k_{B,n},\tau)=\prod_{i=1}^{n}p_{k_{A}|k_{B},\tau}(k_{A,i}|k_{B,i},\tau;d)=\prod_{i=1}^{n}\binom{k_{B,i}}{k_{A,i}}(\tau^{d})^{k_{A,i}}(1-\tau^{d})^{k_{B,i}-k_{A,i}}$
and is maximized at
$\widehat{d}=\frac{\log(\frac{1}{n}\sum_{i=1}^{n}k_{A,i}/\frac{1}{n}\sum_{i=1}^{n}k_{B,i})}{\log(\tau)},$ (1)
which will be denoted as the Binomial ID Estimator (BIDE). Remark that to identify $d$, it is sufficient to fix just two tuning parameters among $t_{A}$, $t_{B}$ and $\tau$. In this paper, we will propose an adaptive data-driven choice for such tuning parameters. It is worth mentioning that the estimator for Discrete Datasets (I3D) presented in Macocco et al., (2023) is obtained setting $\mathcal{M}^{D}\equiv\mathbb{Z}^{D}$ equipped with the $L^{1}$ metric where volumes are computed using the counting measure and Ehrhart polynomials, thus, leading to an alternative non-closed expression of $\widehat{d}$. Moreover, it is also empirically shown by the authors that violating the conditional independence assumption does not induce significant errors as long as $\frac{1}{n}\sum_{i=1}^{n}k_{A,i}\ll n$.
2.2 Largest uniform neighbourhoods in nearest-neighbour methods
We have assumed the existence of a manifold of ID $d$ on which data are (softly) confined, where Euclidean distances measured in $\mathbb{R}^{D}$ are locally well approximated by those measured in an appropriately defined $d$-dimensional manifold tangent space. Accordingly, volumes are taken in $\mathbb{R}^{d}$. Let $r_{i,j}$ be the distance between $i$ and its $j$-th NN with $r_{i,0}$ conventionally set equal to 00, and denote by $B(x_{i},r_{i,k})$ the ball centered at realization $X_{i}=x_{i}$ with radius $r_{i,k}$. The volume of the hyper-spherical shell between neighbours $j-1$ and $j$ of a point $i$ is given by
$v_{i,j}:=\mu(B(x_{i},r_{i,j})\setminus B(x_{i},r_{i,j-1}))=\Omega_{d}(r^{d}_{i,j}-r^{d}_{i,j-1}),$
where $\Omega_{d}=(2\Gamma(3/2))^{d}/\Gamma(d/2+1)$ is the volume of a unit $d$-dimensional hyper-sphere with $\Gamma$ denoting the Euler’s Gamma function. Accordingly, the volume of the $d$-sphere centred in $i$ with radius given by its $k$-th NN is
$V_{i,k}:=\mu(B(x_{i},r_{i,k}))=\sum^{k}_{j=1}v_{i,j}=\Omega_{d}\,r^{d}_{i,k}.$
If we consider the intensity constant and equal to $\rho_{i}$ within a point $i$ and its $k$-th NN at distance $r_{i,k}$, then the hyperspherical shell volumes $v_{i,j}$ with $j\in\{1,\dots,k\}$ are the realizations of $k$ independent and identically distributed (iid) exponential random variables with rate $\rho_{i}$. Therefore, we can write a joint probability density function for $k$ iid exponential random variables $f(v_{i,1},\dots,v_{i,k})=\prod_{j=1}^{k}\rho_{i}e^{-\rho_{i}\,v_{i,j}}$. Thus, looking at the joint density, evaluated at the observed shell volumes $v_{i,1},\dots,v_{i,k}$, as a function of the parameter $\rho_{i}$, and by taking its logarithm we obtain the log-likelihood
${L}_{i,k}(\rho_{i}):={L}(\rho_{i}|v_{i,1},\dots,v_{i,k})=\sum_{j=1}^{k}\left(\log\rho_{i}-\rho_{i}v_{i,j}\right)\,=\,k\log\rho_{i}-\rho_{i}V_{i,k}$
which is maximized at $\widehat{\rho}_{i}=k/V_{i,k}$, matching the $k$-NN local non-normalized density estimator. Its standard deviation, given by $\widehat{\rho}_{i}/\sqrt{k}$, is obtained as the inverse square root of the Fisher information. It is worth noticing the usual bias-variance trade-off: the estimator’s standard deviation decreases as $k$ increases, and, on the other hand, the constant density approximation, on which the exponentiality assumption relies, becomes less solid since the density in a ball of radius $r_{i,k}$ might become non-constant as $k$ increases. Therefore, as we have observed, the $k$-NN density estimator is not immune to the usual bias-variance trade-off problem. However, since the influence of $k$ on the variance can be somewhat decoupled from its impact on the bias, there are margins of improvement. Indeed, to achieve low variance the estimator typically requires a high number of observations or the selection of a large $k$, and for large $k$ a systematic bias is introduced primarily for the statistical units for which the region enclosing $k$ NNs undergoes rapid density variations. In essence, the selected bandwidth for the $k$-NN estimator should be small in regions where the density of data points changes rapidly but it can be larger in areas where the density undergoes slower variations.
Following this reasoning, Rodriguez et al., (2018) developed a quantitative algorithm to find the largest $k$, specific to each data point, that enforces an approximately constant density within a certain confidence level, thus facilitating and automatizing more accurate density estimation in regions where the density exhibits gradual variations. We formalize the procedure to find the optimal $k$ in a statistical hypothesis testing framework. For each point $i$ we consider a growing number of neighbours and iteratively compare two different likelihood models. The first model denoted as $M1$, assumes that the densities of the Poisson process at point $i$ and at its $(k+1)$-th NN are different, say $\rho_{i}$ and $\rho^{\prime}_{i}$. In this case, the maximum log-likelihood computed as a function of $(\rho_{i},\rho^{\prime}_{i})$ is
$L^{M1}_{i,k}(\widehat{\rho}_{i},\widehat{\rho}^{\prime}_{i})=\underset{\rho_{i},\rho^{\prime}_{i}>0}{\max}(L_{i,k}(\rho_{i})+L_{k+1,k}(\rho^{\prime}_{i}))=k\log\Big{(}\frac{k^{2}}{V_{i,k}V_{k+1,k}}\Big{)}-2k$ (2)
The second model denoted as $M2$, on the other hand, assumes no density variation between $i$ and its $k+1$-th NN, leading to
$L^{M2}_{i,k}(\widehat{\rho}_{i},\widehat{\rho}_{i})=\underset{\rho_{i}>0}{\max}(L_{i,k}(\rho_{i})+L_{k+1,k}(\rho_{i}))=2k\log\Big{(}\frac{2k}{V_{i,k}+V_{k+1,k}}\Big{)}-2k.$ (3)
Fundamentally, the volumes $V_{i,k}$ and $V_{k+1,k}$ in the above equations can be taken to be the volumes of hyper-spheres on the intrinsic manifold by appropriately using $d$ instead of the embedding dimension $D$ for their computation. The two maximised log-likelihoods are then compared using a likelihood-ratio test statistic. As long as they are deemed statistically compatible within certain confidence, the putative optimal value for $k$ is increased by one and a new test is carried out; when the likelihood of model ${L}_{i,k}^{M1}$ utilises its extra degree of freedom ($\rho^{\prime}_{i}$) in the maximization to become significantly larger than ${L}_{i,k}^{M2}$ it means that the density variation becomes significant and, thus, the constant density approximation does not hold for the tested $k+1$, meaning that $k$ is the optimal neighbourhood size for point $i$. Formally, by the classical Wilks’ theorem (Wilks,, 1938) the likelihood-ratio test statistic
$\begin{split}D_{i,k}&=-2(L^{M2}_{i,k}(\widehat{\rho}_{i},\widehat{\rho}_{i})-L^{M1}_{i,k}(\widehat{\rho}_{i},\widehat{\rho}^{\prime}_{i}))\\ &=-2k\left(\log V_{i,k}+\log V_{k+1,k}-2\log(V_{i,k}+V_{k+1,k})+\log 4\right)\end{split}$
converges in distribution to a $\chi^{2}_{1}$ random variable, and it can be used to test $H_{0}:\rho_{i}=\rho^{\prime}_{i}$ against $H_{1}:\rho_{i}\neq\rho^{\prime}_{i}$. Since the test can be performed for every $k$ it follows that for each point $i$ we can select $k^{*}_{i}$ as
$k^{*}_{i}=\min\{k:D_{i,k}\geq q_{1-\alpha,1}=:D_{\text{thr}}\}$ (4)
where $q_{1-\alpha,1}$ is the rejection threshold and denotes the $(1-\alpha)$-order quantile of the $\chi_{1}^{2}$ distribution. The choice of $\alpha$ is strongly linked to the usual bias-variance trade-off or the signal-noise disentanglement; in fact, the amount of information that should be regarded as noise is ultimately a choice of the field experts, and the proposed modelling framework allows a clear and transparent control on such choice.
Thanks to this framework, the neighbourhood sizes $k_{1}^{*},\dots,k_{n}^{*}$, are adaptively optimised for each data point, and can be plugged in all the results of the canonical $k$-NN to give birth to an improved $k^{*}$-NN estimator. It is worth noticing that since the average $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}$ certainly scales sublinearly with the number of data points $n$ (unless the data are uniformly distributed), the whole method scales sub-quadratically with $n$. Moreover, to reduce the computational complexity of the method, by fixing the maximum extension of the neighbourhood to a large yet finite value max_$\_k$ (Glielmo et al.,, 2022) the estimation of the optimal $k^{*}_{i}$ is forced to scale linearly with the number of data points $n$, since the likelihood ratio test is performed independently for each point. Some further insights on the scaling of $k^{*}_{i}$ are given in the Appendix.
2.3 Adaptive Binomial Intrinsic Dimension Estimator (ABIDE)
We are now ready to present the new algorithm for ID estimation which automatically uses optimal neighbours. The algorithm can be seen as an extension and improvement of BIDE. In particular, the idea revolves around the concept of exploiting the largest possible neighbourhoods for the computation of the ID while, at the same time, satisfying the only required assumption, namely the local constant density hypothesis encapsulated in the Poisson process at the basis of the estimator. Concretely, this means computing the ID with the binomial estimator fixing the size of the neighbourhood not to a constant value $k$ but to the adaptively chosen values $k^{*}_{i}$ found by the optimisation procedure described above. In this manner we can select, separately for each data point $X_{i}$, the largest possible neighbourhood where the assumption of constant density is not rejected. The distance of the $k_{i}^{*}$-th neighbour from point $i$ is then $r_{i,k_{i}^{*}}$.
There is an important caveat that makes the methodology much more challenging. The algorithm that returns the optimal $k_{i}^{*}$ requires the knowledge of the ID to compute volumes of hyper-spheres in the proper intrinsic manifold (see (2) and (3)). To address this key issue we propose the following iterative procedure. An initial value of ID is found using a standard method like the BIDE or the Two Nearest Neighbours (2NN) estimator Facco et al., (2017). Since the 2NN is well-established, it does not have tuning parameters and it focuses on very small neighbourhoods of order 2, we adopt it as the starting value for $d$. Moreover, it is interesting to note that the 2NN in its maximum likelihood formulation is equivalent to BIDE under certain conditions. To see this, we recall that the 2NN estimator in its (unbiased) maximum likelihood formulation (Denti et al.,, 2022) is defined as
$\widehat{d}_{2NN}=\frac{n-1}{\sum_{i=1}^{n}\log(r_{i,2}/r_{i,1})},$
therefore, if we allow the ratio $\tau$ in (1) to be point specific and, we set it as $\tau_{i}=r_{i,1}/r_{i,2}$ and choose the radii such that $\sum_{i=1}^{n}k_{A,i}/\sum_{i=1}^{n}k_{B,i}=1/e$ we obtain
$\widehat{d}=\frac{\log(1/e)}{\frac{1}{n}\sum_{i=1}^{n}\log(\tau_{i})}=\frac{n}{\sum_{i=1}^{n}\log(r_{i,2}/r_{i,1})}\approx\widehat{d}_{2NN}.$
Thus, fixing the count ratio and allowing $\tau$ to be point specific we have that the BIDE estimator is equivalent to the 2NN estimator in its basic maximum likelihood formulation, i.e. with no bias correction. Moreover, for the same fixed count ratio a simple calculation, setting $\widehat{d}=n/\sum_{i=1}^{n}\log(r_{i,2}/r_{i,1})$ and solving for $\tau$, shows that to get the equivalence of the two estimators we should have $\tau=\exp(n^{-1}\sum_{i=1}^{n}\log(r_{i,1}/r_{i,2}))$ which is the geometric mean of $\{r_{i,1}/r_{i,2}\}_{i=1,\dots,n}$.
Now, proceeding with the description of our iterative procedure, once we have set a starting value for $d$, the optimal set of $k^{*}_{1},\dots,k^{*}_{n}$ is then computed according to (4), and used as input for BIDE which gives a new ID estimate. Such ID, typically different from the one obtained with the 2NN, is used to compute another set of $k^{*}_{1},\dots,k^{*}_{n}$, that will become the input for an updated ID estimate. The procedure continues until convergence, which occurs simultaneously on both the ID and the $k_{1}^{*},\dots,k_{n}^{*}$.
More precisely, at each iteration, once $k_{1}^{*},\dots,k_{n}^{*}$ have been computed, $d$ is estimated using BIDE as in (1): it suffices to set $k_{B,i}=k_{i}^{*}-1=:k_{B,i}^{*}$ and $t_{B,i}(k_{i}^{*})=r_{i,k_{i}^{*}}$ as the radius of the open ball containing $k_{B,i}^{*}+1$ points. Successively, we compute $k_{A,i}^{*}$ as the number of points contained in the open ball of radius $t_{A,i}(k_{i}^{*})=t_{B,i}(k_{i}^{*})\tau$ without counting point $i$, and we set $k_{A,i}=k_{A,i}^{*}$. Furthermore, we can devise an optimal procedure to select the ratio $\tau$ of the radii of the two open balls. As shown in the supplementary material of Macocco et al., (2023), minimizing the asymptotic variance of the estimator in (1) leads to the optimal choice $\tau\approx 0.2032^{1/d}$. Since the optimal $\tau$ depends on $d$, the true ID, we can naturally include the computation of $\tau$ in the iterative procedure by using the latest estimate of the ID at every step in place of its true value. We refer to this procedure as Adaptive Binomial ID Estimator (ABIDE) and denote the resulting estimator with $d^{*}$.
Figure 1: The ABIDE algorithm, thanks to its iterative nature, progressively adjusts the estimates of the ID and the size of (approximately constant density) neighbourhoods, allowing it to escape the noise scale and to find the true ID of the underlying manifold. The first row shows how, throughout the iterations, the size of the neighbourhoods for two selected points $i$ and $j$ grows as the ID estimates lower from 2 down to 1. In the second row, we report the evolution of the ID estimate and the boxplots representing the subsequent distributions of the $k^{*}$.
In Figure 1 we report a sketch of our algorithm on a toy dataset, created in the following way. We consider two one-dimensional Gaussian distributions centred, respectively, in $\frac{\pi}{2}$ and $\frac{5}{3}\pi$, and with standard deviations of 1 and 0.5, and we sample $X_{1},\dots,X_{1000}$ points, 500 from each one of the two. We then map the points on a curved manifold by adding a second coordinate given by $Y_{i}=\sin(X_{i})$. Finally, we add a Gaussian noise of standard deviation 0.025 along the second coordinate $Y$, which sets the scale below which the dataset appears to be 2-dimensional. The bottom left panel shows how the one-dimensional nature of the data emerges throughout the successive iterations of ABIDE. Simultaneously, the neighbourhoods where the density can be considered approximately uniform get larger and larger as shown in the bottom right panel. The reason is that, as long as the estimated ID is approximately 2 (which is the value obtained using 2NN at iteration 0) the manifold is expected to be 2-dimensional, and thus the uniform density test is performed onto 2-dimensional hyper-spheres (disks). Accordingly, the size of the uniform neighbours is limited to scales comparable with the amplitude of the noise along the $Y$ coordinates, beyond which the density of data points belonging to a (supposedly) 2-dimensional manifold drops to 0. However, the extension of such neighbourhoods is large enough so that, at iteration 1, the estimated ID using BIDE in combination with the $k_{i}^{*}$ is of order 1.3. Consequently, as the estimated ID gets closer to one, the uniform density test is performed on a 1-dimensional manifold, allowing to obtain larger values for $k^{*}$, since the neighbours’ size can extend along the proper manifold, well beyond the scale of the noise.
This example makes the interplay between the estimated ID and the size of uniform neighbours evident: it is exactly because of larger neighbourhoods that it is possible to escape the scale at which the noise is relevant and, thus, uncover the true dimensionality of the data. At the same time, the proper size of the neighbourhoods is found using the right ID. From this perspective, it is evident how the process is intrinsically iterative.
As we will show in the Results section, in most cases only a few iterations are necessary to get a stabilization of the ID estimate and of the associated $k^{*}_{i}$ values. Intuitively, this can be understood by the fact that the ratio $\sum_{i=1}^{n}k^{*}_{A,i}/\sum_{i=1}^{n}k^{*}_{B,i}$ is not strongly affected by a possibly inaccurate initial estimate of the ID, since both numerator and denominator grow proportionally if the data generating density is sufficiently regular. In the Appendix, we provide a formal justification for the fast convergence of Algorithm 1. Specifically, we show that Algorithm 1 is a fixed point iteration and use some theoretical results regarding the large sample behaviour of the associated contraction mapping to show that, for $n$ sufficiently large, very few iterations are needed to achieve numerical convergence at a given tolerance $\delta$. In practice, we observe that 5 iterations are always enough to achieve a stabilization of the estimates with a tolerance $\delta\asymp 10^{-4}$, and most of the improvement is obtained in the first 2 iterations. Thus, the suggested iterative procedure is highly scalable and brings substantial improvement over the starting estimate obtained using the 2NN with little additional computational cost. In the numerical experiments reported below, instead of selecting a value for $\delta$, we choose to fix the number of iterations to 5. ABIDE is formally summarised in Algorithm 1. In the Appendix, we provide theoretical guarantees regarding the numerical convergence of Algorithm 1 and, we show consistency and asymptotic normality of the resulting ABIDE estimator. Moreover, we show that for large sample sizes the local homogeneity assumption of the Poisson process and the independence assumption of the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$ tend to be asymptotically fulfilled under rather flexible conditions on the intensity function $h$ of the underlying Poisson process.
Algorithm 1 Adaptive-BIDE workflow
1:$d_{\text{current}}\leftarrow$ ID obtained from the 2NN estimator
2:$d_{\text{next}}\leftarrow 0$
3:for  $it<$ max_iter  do
4:     $\tau=0.2032^{1/d_{\text{current}}}$
5:     for  $i<n$  do
6:         compute $k^{*}_{i}$ (using $d_{\text{current}}$) and set $k_{B,i}^{*}=k_{i}^{*}-1$
7:         $t_{B,i}(k_{i}^{*})=r_{ik_{i}^{*}}$
8:         $t_{A,i}(k_{i}^{*})=\tau\,t_{B,i}(k_{i}^{*})$
9:         $k_{A,i}^{*}=\sum_{j=1}^{n-1}\mathbf{1}\{t_{A,i}(k_{i}^{*})-r_{i,j}>0\}$
10:     end for
11:     $d_{\text{next}}=\frac{\log(\frac{1}{n}\sum_{i=1}^{n}{k_{A,i}^{*}}/\frac{1}{n}\sum_{i=1}^{n}{k_{B,i}^{*}})}{\log(\tau)}$
12:     if $|d_{\text{current}}-d_{\text{next}}|<\delta$ then
13:         break
14:     end if
15:     $d_{\text{current}}=d_{\text{next}}$
16:end for
17:$d^{*}=d_{\text{next}}$
18:for  $i<$ n  do
19:     compute $k^{*}_{i}$ (using $d^{*}$)
20:end for
21:return $d^{*},\;k^{*}_{i}$
The asymptotic normality of ABIDE (also derived in the Appendix) allows for a simple uncertainty quantification of the estimates bypassing the need to use computationally intensive resampling procedures such as bootstrap or jackknife. Indeed, the approximate (asymptotic) confidence interval reads
$\Big{[}d^{*}-\frac{z_{1-\beta/2}}{\sqrt{nI(d^{*})}},d^{*}+\frac{z_{1-\beta/2}}{\sqrt{nI(d^{*})}}\Big{]}$
where $z_{1-\beta/2}$ is the quantile of order $21-\beta/$ 2 of the standard Gaussian distribution and $I(d^{*})$ is the observed Fisher information evaluated at $d^{*}$ which is given by
$I(d^{*})=\frac{\log(\tau)^{2}\tau^{d^{*}}n^{-1}\sum_{i=1}^{n}k_{B,i}^{*}}{1-\tau^{d^{*}}}.$ (5)
Note that the data-driven optimal choice $\tau=0.2032^{1/d}$ implies that $\operatorname{Var}[d^{*}]\asymp d^{2}/nk^{*}_{B}$. Moreover, we do not consider additional inflating terms in the asymptotic variance due to possible dependencies among the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$, as we show in the Appendix, both theoretically and numerically, that these are negligible for large sample size.
To assess the reliability of the estimated ID and the selected neighbourhoods, one can exploit a similar model validation framework proposed in Macocco et al., (2023). In particular, at each iteration and for the values obtained at convergence, the set of $k^{*}_{i}$ and the estimated ID are used to build a theoretical distribution with pmf $\widetilde{p}_{k_{A}^{*}}$, obtained as a mixture of binomial distributions with pmf $p_{k_{A}^{*}|k_{B}^{*}}$ and mixing pmf $p_{k_{B}^{*}}$ given by the empirical pmf of $k_{B,1}^{*},\dots,k_{B,n}^{*}$. Thus, letting $p_{k_{A}^{*}}$ be the empirical pmf of the observed $k_{A,1}^{*},\dots,k_{A,n}^{*}$, the theoretical mixture distribution with pmf
$\widetilde{p}_{k_{A}^{*}}(x)=\sum_{y\geq 0}p_{k_{B}^{*}}(y)p_{k_{A}^{*}|k_{B}^{*}}(x|y)$
is compared to $p_{k_{A}^{*}}$ through a two-sample test for equality in distribution, which is performed between an arbitrarily large artificial sample drawn from $\widetilde{p}_{k_{A}^{*}}$ and the observed $k_{A,1}^{*},\dots,k_{A,n}^{*}$. Differently from Macocco et al., (2023), who adopt a standard Kolmogorov-Smirnov test, in our analysis we choose the Epps-Singleton test (Epps and Singleton,, 1986), which is based on the empirical characteristic function and is also valid for discrete distributions. In the goodness-of-fit analysis, a large p-value of the test indicates that the modelling assumptions are more likely to be met and therefore the ID estimate is reliable.
3 Numerical experiments
In this section, we illustrate the behaviour of ABIDE and comment on its performances, in particular against non-iterative algorithms. Assuming that in the applied analyses we want to achieve robustness to low or moderate noise magnitudes, we opt for a conservative approach regarding the null hypothesis of constant density, enforcing a moderately high sensitivity to density variations. This means that it must be relatively easy to reject $H_{0}$, and thus accept the alternative hypothesis that states that the density is effectively changing. To this aim we set $\alpha$ to the standard value of 0.01 and, accordingly, take $D_{\text{thr}}=6.635$. This is a cautionary choice, as it ends up selecting smaller neighbourhoods. Smaller values of $\alpha$ would result in the selection of larger neighbourhoods, smaller variance, and higher robustness of the estimator to noise, at the cost of potentially introducing some bias in non-noisy scenarios. A numerical experiment shows the effect of this choice in the Appendix, and the relevance of this trade-off will be the object of future investigations.
3.1 Noisy Gaussian distribution
The first simulated scenario consists of a $d$-dimensional multivariate standard Gaussian random variables with independent components embedded in a $D$-dimensional space. The embedding is obtained by adding noise to all the $D$ coordinates. More precisely, let us consider the random vector $(S_{1},\dots,S_{d})\sim\mathcal{MN}(0_{d},\Sigma)$. Let $\Sigma=\sigma_{S}I_{d}$, and consider $\varepsilon_{1},\dots,\varepsilon_{D}$ independent copies of $\epsilon\sim\mathcal{N}(0,\sigma_{\varepsilon})$ where $\sigma_{S}$ and $\sigma_{\varepsilon}$ represent respectively the signal and the noise level and thus $\sigma_{\varepsilon}/\sigma_{S}$ is the noise-to-signal ratio. In the notation used above, we let $X_{i}=(S_{1}+\varepsilon_{1},S_{2}+\varepsilon_{2},\dots,S_{d}+\varepsilon_{d},\varepsilon_{d+1},\dots,\varepsilon_{D})_{i}$ and consider independent copies $X_{1},\dots,X_{n}$ representing the rows of the data matrix. In the simulation study, we compare ABIDE with the 2NN estimator which is also adopted as the starting point of the ABIDE iteration. To assess the variability of the estimators we generate 500 independent Monte Carlo replicas with sample size $n=5,000$, and, compute the Monte Carlo empirical quantiles to construct the Monte Carlo 99% confidence interval. We consider $\sigma_{S}=1$ and let $\sigma_{\varepsilon}$ range in $[10^{-4},10^{-1}]$. For all the experiments we set $D=100$ and $d\in\{2,5\}$. This is a misspecified setting because the ID that we aim to retrieve is the one that we would obtain if the data were non-noisy. More precisely, misspecification means that there is a mismatch between the $d$ we aim to estimate and the usual ID definitions that lead to the construction of estimators. Indeed, it is not possible to project the highly noisy data into a $d$-dimensional space without loss of information. Moreover, the dimension of the support of the probability distribution generating the data turns out to be larger than $d$, and this is the typical case when dealing with real data. Nevertheless, Figure 2 shows that the proposed estimator behaves well in terms of bias and variance, and, compared to the 2NN, improves protection against noisy misspecified scenarios and achieves smaller uncertainty (shaded areas) on the estimates.
Figure 2: Estimated ID as a function of embedding noise for two Gaussian datasets of different $d$ and with $n=5,000$ points. The shaded area is the Monte Carlo 99% confidence interval and the black dotted line is the true $d$.
3.2 Noisy-curved-inhomogeneous 2-dimensional distribution
The second dataset we analyse is adapted from d’Errico et al., (2021). Its original form, which we reported in Panel A of Figure 3, shows a uniform background with eight Gaussian distributions with different means and variances. As it is, such a dataset can be analysed with canonical tools, as any estimator would find an ID close to 2. To make the experiment more interesting, we first embed such points on a 3-dimensional Möbius strip, which is then placed within a 20-dimensional space by adding a 20-dimensional Gaussian noise of scale $\sigma_{\varepsilon}=10^{-3}$.
Figure 3: Panel A: the original 2-dimensional distribution of the 20,000 artificial data points under examination (taken from d’Errico et al.,, 2021). The description of the embedding procedure to obtain the actual dataset is described in the main text. Panel B: ID estimation using the standard non-adaptive Binomial Intrinsic Dimension Estimator (BIDE) at fixed radius or scale $t_{B}$ (blue circles) with associated p-values obtained through model validation (red squares). The isolated starred points represent, respectively, the ID estimate (blue) and the p-value (red) obtained with the ABIDE estimator. As a reference, the latter is placed in the average (over all data points) of the distance of the $k_{i}^{*}$ neighbour: $\frac{1}{n}\sum_{i=1}^{n}{t_{B,i}(k^{*}_{i})}$. The dashed line (reported also in other panels) is the theoretical ID=2 of the original dataset before adding noise. Details and considerations in the main text. Panel C: ID estimation and p-values using BIDE at fixed neighbourhood size $k$ (on the x-axis). Also in this case ABIDE results are reported as the starred values, in correspondence of $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}$. Panel D: points are coloured according to their $k^{*}_{i}$ value to visualize the concept of adaptive neighbourhood. Panel E: the evolution of the ID (blue) and the associated p-values (red) from model validation at the successive iteration of ABIDE. The final values at convergence are starred and reported also in the BIDE plots for comparison. Panel F: 2NN vs ABIDE and p-values as a function of the amplitude of the noise used to embed the data.
As a first test, we try to understand whether it is possible to observe a plateau at ID=2 as prescribed by conventional estimators by tuning the fixed radius, $t_{B}$, in BIDE. To check the goodness of the estimate we also perform the model goodness-of-fit test through the Epps-Singleton test and look at the obtained p-values. The results are reported in Panel B of Figure 3 and we cannot find the plateau around the ground truth ID value of 2. Instead, the estimated ID keeps decreasing as $t_{B}$ increases. However, the presence of non-negligible noise, still relevant at such small scales, does not allow us to find the ID of the manifold within such neighbourhoods, as those are too small to overcome the noise amplitude. Conversely, as soon as one measures the ID at larger scales by increasing $t_{B}$ to escape the noise, many of the selected neighbourhoods undergo high-density variations. Accordingly, the p-values start declining together with the ID. As a consequence, if one uses the goodness-of-fit of the model as a criterion for choosing the radius $t_{B}$, one would be forced to conclude that the ID of this data set is significantly larger than the ground truth value.
To try to account for density variations, we fixed the neighbourhood size and computed the ID using BIDE with $k$ fixed. Also in this case the results, reported in Panel C of Figure 3, are unsatisfactory. Even if the ID appears to slowly converge to a value close to 2 as $k$ grows, the relative p-values are extremely low, which would lead to consider the ID obtained with a large $k$ unreliable. Density variations are too strong and this approach does not allow us to make reliable predictions, even if the ID trend is going in the proper direction.
Next, we apply the ABIDE algorithm. As described above, the algorithm does not require fixing either $t_{B}$ or $k$, as those values are chosen automatically. The result of the ABIDE estimator is given by the ID (and its p-value) and the $k^{*}$ distribution (and eventually $t_{B}(k^{*})$) at convergence (starred values in Panel E of Figure 3). Therefore we do not have to change any parameter explicitly to find patterns (like plateaus) in the ID path. The obtained ID and p-values are reported as starred isolated points in Panels B and C of Figure 3 in correspondence of, respectively, $\frac{1}{n}\sum_{i=1}^{n}{t_{B,i}(k^{*}_{i})}$ and $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}$. Quantitatively speaking, we obtained ID $\approx 2.35$, close to the true value of 2, and a p-value $\approx 5\times 10^{-7}$, which is many orders of magnitudes larger than the ones obtained at the same (averaged) scale or at fixed (for every point) neighbourhood size. This is a strong indication that choosing the neighbourhood size in a point-wise adaptive manner is the key approach if we want to obtain a statistically reliable description of the data distribution.
We now analyse in more detail how the ABIDE results are obtained. Panel E of Figure 3 shows the convergence of the ID (blue) to a value slightly above 2 within three iterations (the ID at iteration 0 was obtained employing the 2NN estimator). The corresponding p-values (red) are always of order $10^{-7}$ or $10^{-6}$, meaning that thanks to the adaptive neighbourhood implementation, we are properly exploring regions with a fairly constant density around each datapoint, resulting in a proper binomiality of the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$. The values at convergence are then considered the results of the ABIDE estimator. These have been starred and reported, allowing for a direct comparison, in Panels B and C.
In the Appendix, we also report the distributions of $k_{i}^{*}$ and the related radii $t_{B,i}(k^{*})$ obtained with 2NN and ABIDE. As occurred in the illustrative example of Figure 1, the lowering of the ID is paired with an enlargement of neighbourhoods’ size. To visualize the concept of adaptive neighbourhood, in Panel D we coloured the points of the dataset according to their value of $k^{*}$ obtained with ABIDE (for visualization reasons we used the original 2-dimensional dataset). As one can see, points close to the border of the distribution present a small $k^{*}_{i}$, as the density sharply changes within a small number of neighbours. Conversely, points at the centre of a density peak show a much larger value of $k^{*}_{i}$, as the density varies in a sufficiently smoother way to be considered approximately constant on such a scale.
As a last experiment, we consider the same dataset for different noise intensities. In particular, we use a noise of $10^{-4}\leq\sigma_{\varepsilon}\leq 10^{-2}$ to embed the original 2-dimensional data points. We expect that the stronger the noise the harder it is to grasp the true dimensionality of the hidden manifold. A high noise intensity necessarily implies an overestimate of the manifold dimension, as more orthogonal coordinates are required to represent the data. Panel F of Figure 3 reports the ID estimates obtained by 2NN and ABIDE (and its associated p-values) as a function of the noise intensity. We observe how ABIDE is capable of finding the ID of the data manifold much more robustly than the 2NN method, being able to tolerate a higher level of noise remaining much closer to the ground-truth ID value. Indeed, as far as $\sigma_{\varepsilon}\lesssim 10^{-3}$, ABIDE correctly finds ID estimates very close to 2 and, correspondingly, p-values of order $10^{-1}$. Differently, when $\sigma_{\epsilon}\gtrsim 10^{-3}$, the noise becomes so intertwined with the manifold structure that not even ABIDE is capable of getting reasonable estimates of ID. Accordingly, the associated p-values drop dramatically, signalling that the estimates are no longer quantitatively reliable in such a regime.
Already within this artificial example, we can identify the many advantages offered by ABIDE. Thanks to its iterative formulation, it can sequentially adjust the ID and the neighbourhood sizes to point-wise identify the largest region where the density is approximately constant and enforce the hypotheses behind the estimator. At the same time, such a procedure allows us to surpass the barrier created by noise with higher capability than other methods and identify a proper adaptive scale at which the ID should be computed.
3.3 Applications to image data and molecular dynamics trajectories
In this section, we present the results of the ABIDE estimator on three real-world datasets with the following specifications. The first one is the so-called OptDigits (Alpaydin and Kaynak,, 1998), a collection of data matrices of dimension $n=3,823$ and $D=64$, representing small-sized hand-written digits. For the second application, we remain in the realm of hand-written digits with the paradigmatic MNIST dataset (LeCun et al.,, 2010), with size $n=70,000$ and $D=784$. Being an order of magnitude larger than the previous dataset concerning both the number of entries and features, it represents a very different challenge despite the similar nature of what is encoded in the data. As a last example, we analyse a biomolecular trajectory. We use $n=3,758$ uncorrelated frames from a replica-exchange molecular dynamics simulation (Sugita and Okamoto,, 1999) at a temperature of 340 K of the CLN025, a small peptide made up of 10 residues that folds into a beta-hairpin (Honda et al.,, 2004). Each frame of the trajectory, initially computed in (Carli and Laio,, 2021) and corresponding to a single data point, can be represented in many different ways. We decided to work with the compact but still meaningful representation given by the dihedrals angles (Bonomi et al.,, 2009; Cossio et al.,, 2011). With this description, the dimensionality of the dataset is $D=32$. Since we are dealing with angles, the periodicity of $2\pi$ has to be taken into account when computing the distances between data points.
Figures 4 and  5 show the analyses carried out similarly to what we did in the previous section on artificial data. In particular, for each one of these figures, the first row shows some samples extracted from the datasets. In the second row, from left to right, we report: ID (blue) and p-values (red) estimated with ABIDE at successive iterations; BIDE ID scaling at fixed $t_{B}$ with ABIDE results superimposed (as a starred blue bullet for the ID and a starred red square for the p-value); BIDE ID scaling at fixed $k$ with ABIDE results superimposed (as starred bullets); the distribution of $k^{*}$ computed using the 2NN estimate and the ABIDE estimate.
Figure 4: In the two upper rows, we report ABIDE performances on the OptDigits dataset made up of $n=3,823$ elements with embedding dimension $D=64$, together with a sample of 6 data points (first row). The first panel shows the evolution of the ID (blue), and of the p-value (red), with successive ABIDE iterations. In particular, one can appreciate how the ID estimated with ABIDE converges within 3 iterations and is significantly smaller than the 2NN estimate (from 10.43 to 8.05). The second and third panels show the scaling of the ID obtained with BIDE by fixing the radius $t_{B}$ or by fixing the number of neighbours $k$. Since no plateau is present in the scaling curves, without ABIDE we would not have a proper criterion to recognize a meaningful ID. Conversely, ABIDE - shown as starred points centred, respectively, in $\frac{1}{n}\sum_{i=1}^{n}{t_{B,i}(k^{*}_{i})}\approx 3.5$ and $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}\approx 9$ allows us to uniquely identify the point-dependent scale as the largest neighbourhood size $k^{*}_{i}$ where the local density is approximately constant. The last panel shows the distribution of $k^{*}$ computed with the ID estimated by 2NN and with ABIDE at convergence.
In the two bottom rows, we report ABIDE performances on the MNIST dataset, with sizes $n=70,000$ and $D=784$, together with a sample of 6 data points (top row). In the first panel, we observe how the ID drops from the initial value of 15.19 to 13.47. The second and third panels show the ID estimate at a fixed radius $t_{B}$ and at a fixed number of neighbours $k$, respectively. We notice that no plateaus are present and that p-values quickly become very low (in the case of fixed radius they are numerically 0). Conversely, ABIDE shows how the adaptive neighbourhood size allows for an ID estimate with a significantly larger p-value at an average scale $\frac{1}{n}\sum_{i=1}^{n}{t_{B,i}(k^{*}_{i})}\approx 16$, corresponding to an average neighbourhood size of $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}\approx 6.5$. Again, in the rightmost panel, one can observe the distributions of $k^{*}_{i}$ with 2NN and ABIDE.
We begin by highlighting the common features of all three analyses. We observe that the ID always decreases from the initial 2NN estimate (Iteration 0 in the first panels). The reason is most likely related to the high sensitivity of 2NN to the noise, which unavoidably leads to overestimating the ID.The same phenomenon occurs for BIDE estimates at small, fixed radii (second panels): even if the density is likely to be locally constant and, accordingly, the p-values are typically high, this is the regime where noise plays a relevant, if not dominant, role in “smearing” and hiding the true manifold ID. Unfortunately, when exploring the ID at larger $t_{B}$ values, that allow us to overcome the noise, we observe that the p-value drops consistently, meaning that for many data points the selected neighbourhood has significant density variations.
A similar behaviour occurs when the neighbour’s size $k$ is fixed (third panels): for small $k$ the estimates are potentially reliable but are affected by noise; conversely, when the neighbourhood becomes larger, the density is less likely to be uniform. In all cases we indeed observe a broad distribution of $k^{*}$ (rightmost panel), suggesting that, for this type of data, the ID estimate at fixed neighbourhood size (or even worse, at fixed radius) is inevitably poor. Last, but not least importantly, we also observe the absence of a clear plateau in each of the BIDE scaling plots (second and third panels of each figure). This means that it is typically not possible to unequivocally decide which is the correct ID and the corresponding scale at which it is meaningful to observe the system. Using ABIDE we are instead capable of automatically identifying a precise scale, where the ID estimate is close to the true one and is generally associated with a relatively high p-value. In the same fashion as what we did for the artificial dataset, we report ABIDE results as starred points in the ID/p-value vs $t_{B}$ and ID/p-value vs $k$ panels obtained with the canonical scaling algorithms of the BIDE estimator.
In particular, for the OptDigits we obtain ID=8.05 (with a p-value $\approx 10^{-14}$), starting from an ID=10.43 computed with 2NN. Concerning the neighbourhoods, we register $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}\approx 9$. For MNIST we have an ID decreasing from 15.19 to 13.47 (with p-value $\approx 10^{-55}$) and $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}\approx 6.5$. Finally, for the peptide, we have an ID converging at 5.55 (p-value $\approx 10^{-25}$) starting from an initial value of 6.78 and $\frac{1}{n}\sum_{i=1}^{n}{k^{*}_{i}}\approx 10$.
Figure 5: ABIDE performances on the dihedral representation of CLN025 peptide, a dataset of size $n=3,758$, $D=32$. The first row shows three typical configurations of the peptide (from left to right): folded beta-hairpin, misfolded twisted, and unfolded. From the first panel of the second row, we can appreciate how the ID drops by 18%, from the 2NN value of 6.78 to the ABIDE one of 5.55. BIDE estimates obtained by selecting the average ABIDE scale are close to the ones obtained with ABIDE (second and third panels).
4 Discussion
We introduced an approach with provable theoretical guarantees aimed at mitigating some problems commonly affecting ID estimators: the choice of the scale at which the ID should be measured and the dependence of the ID estimate on the noise. These two issues are strongly intertwined with the verification/enforcement of the statistical hypotheses on which the ID estimate is based. Starting from a likelihood formulation of the $k$-NN method, we leveraged an algorithm that finds, for each data point, the largest neighbours where the density is approximately constant (at a given confidence level). This, we have shown, allows us to enhance the statistical accuracy of the ID estimate.
To prove the effectiveness of this approach, which we called ABIDE, we tested it on both artificial and real-world datasets. The former, in which we had full control and comprehension of the data, allowed us to understand how ABIDE works and have an idea of its capabilities. In particular, we addressed its robustness against noise and verified the reliability of its estimates, exploiting the model validation developed in Macocco et al., (2023). We then tackled different real-world datasets and, in each one of them, ABIDE managed to deliver insights that could not be obtained with other methods. In particular, in all of the cases we considered, the ID does depend on the scale. In these conditions, strictly speaking, the question “What is the ID of this dataset?” cannot be answered unequivocally. Using ABIDE, we can say that a meaningful scale at which the ID should be measured is the largest scale at which the density is constant. Clearly, in the presence of noise, the density will be constant even at lower scales, but the ID measured at such scales will be spuriously increased by the noise itself.
The results of the approach depend on a single hyper-parameter with a clear interpretation, the maximum probability of rejecting the hypothesis of constant density ($\alpha$ in (4)). We plan to systematically study how the results are affected by this parameter. In the analyses performed we used $\alpha=0.01$ corresponding to a statistical confidence of 99% of correctly rejecting the constant density hypothesis when performing the test. On the one hand, being more tolerant of noise allows for enlarged neighbourhoods. At the same time, using smaller values of $\alpha$ one is less stringent on the requirement of density’s smoothness. To check the potential of the approach introduced with ABIDE, it would also be interesting and important to test the same framework on other suited ID estimators, like GRIDE (Denti et al.,, 2022) whose extension is briefly outlined in the Appendix, and the maximum likelihood estimator proposed in Levina and Bickel, (2004), just to mention two of them where the method we propose could be straightforwardly extended. The possibility of finding proper and meaningful neighbourhood sizes can be of great help in many relevant fields of data analysis. In particular, any method relying on $k$-NN statistics would benefit from the algorithm we have introduced in this paper. Our approach naturally extends to virtually any tool where the choice of the nearest-neighbour order to be considered is left to the user as a tuning parameter without being principle-based. Among such methods, we mention density estimation techniques, where the adaptive neighbourhood was conceived (Rodriguez et al.,, 2018) and has already found some applications (Carli and Laio,, 2021; Carli et al.,, 2024); clustering routines, both those density-based (d’Errico et al.,, 2021) and neighbours-graph-based (Tobin and Zhang,, 2024); projective methods like Isomap (Tenenbaum et al.,, 2000), UMAP (McInnes et al.,, 2018) and Locally Linear Embedding (Roweis and Saul,, 2000). For all the aforementioned methods, the sole selection of an adaptive neighbourhood size might improve their performances. Such developments are left to future research. Finally, after we have grasped the capabilities of the ABIDE estimator in Euclidean spaces, it might be interesting to extend it to discrete metric spaces.
Appendix
The following sections will be devoted to the derivation of theoretical guarantees for the ABIDE estimator introduced in the paper. Moreover, we briefly show how the methodology can be extended to its Bayesian counterpart. As a proof of concept, we show that GRIDE, another ID estimator recently proposed by Denti et al., (2022), can be improved by selecting optimal neighbourhoods of data points. In principle, this could be done for other ID estimators based on NNs.
Appendix A Theoretical guarantees
We provide theoretical guarantees regarding the numerical convergence of the proposed algorithm (Proposition 1) and show consistency of the resulting ABIDE estimator (Proposition 2). Moreover, we also prove asymptotic normality (Proposition 3). Throughout the section we will use the following notation:
Notation.
In the following, we will denote constants (sequentially) by $c_{1},c_{2},c_{3},\dots$ and so on. Given two functions $f_{n}$, $g_{n}$ and two positive universal constants $c_{1}\geq c_{2}$, we will use the notations $f_{n}\sim g_{n}$, $f_{n}\lesssim g_{n}$, $f_{n}\gtrsim g_{n}$ and $f_{n}\asymp g_{n}$ to indicate respectively: $f_{n}/g_{n}\to 1$ (as $n\to\infty$) and $f_{n}\leq c_{1}g_{n}$, $f_{n}\geq c_{2}g_{n}$, $c_{2}g_{n}\leq f_{n}\leq c_{1}g_{n}$ (for all $n$ sufficiently large). Probabilistic statements will be implicitly made with respect to the usual product probability space $(\Omega^{(n)},\mathcal{A}^{(n)},P^{(n)})$ on which the collection of random vectors $X_{1},\dots,X_{n}$ is defined, and such statements will be often made as $n\to\infty$. For notation simplicity, we omit the superscript in the probability function so that $P:=P^{(n)}$. For a certain event, we will abbreviate almost surely by a.s. to indicate that it happens with probability 1. For a certain property, we will abbreviate almost everywhere by a.e. to indicate that it holds everywhere up to a set of measure zero. For asymptotic results regarding ABIDE, we will subscript the estimator $d^{*}=:d_{n}^{*}$.
Before proceeding it is also necessary to point out that, for extremely high-dimensional data, methods based on NN are affected by an underestimation bias which may be minimized through our procedure, however, it cannot be reduced to zero. This is a consequence of the fact that a very large ID causes density variations to occur at very small scales. In the following theoretical results, for the sake of simplicity, we implicitly assume that the density of the data, and equivalently the intensity function $h$ of the Poisson process, satisfies some regularity conditions to make such structural bias asymptotically negligible:
Assumption 1.
The intensity function $h$ of the underlying Poisson process (introduced in Section 2) is bounded, Lebesgue integrable and continuous a.e.
A.1 Supporting lemmas
We start with the following lemmas.
Lemma 1.
As $n\to\infty$, $k_{i}^{*}/n\to 0$ for all $i=1,\dots,n$.
Proof.
We employ the results on order statistics and NN distances reported in Chapter 1, Section 1.2 of Biau and Devroye, (2015). For every $i=1,\dots,n$, we condition on point $i$ and its $k+1$-th NN obtaining $V_{i,k}\sim V_{k+1,k}\sim U_{(k)}$, where $U_{(k)}$ is the $k$-th order statistic of a iid uniform random vector $(U_{1},\dots,U_{n-2})$. Moreover, $U_{(k)}\sim\mathrm{Beta}(k,n-1-k)$ which implies that
$V_{i,k}\sim V_{k+1,k}\asymp\frac{k}{n},$
with probability tending to 1. Thus, since $k_{i}^{*}$ is defined as
$k_{i}^{*}=\sum_{k=1}^{n-2}k\mathbf{1}\{D_{i,k-1}<q_{1-\alpha,1}\cap D_{i,k}\geq q_{1-\alpha,1}\},$
it suffices to study the behaviour of the test statistic $D_{i,k}$. By definition of $D_{i,k}$ we obtain
$D_{i,k}$ $=-2k\left(\log V_{i,k}+\log V_{k+1,k}-2\log(V_{i,k}+V_{k+1,k})+\log 4\right)$
$\asymp-2k(2\log(k/n)-2\log(k/n)+\log(4))$
$=k\log(1/16)=O_{P}(1)$
for $n$ sufficiently large, and the thesis follows since $k_{i}^{*}=O_{P}(1)$ for all $i=1,\dots,n$. ∎
Lemma 2.
The distribution of the random vector $X$ is locally uniform on the set of balls $\{B(x_{i},r_{i,k_{B,i}^{*}})\}_{i=1,\dots,n}$ as $n\to\infty$.
Proof.
First, consider a compact set $K\subset\mathbb{R}^{D}$ and partition it into $S$ subsets denoted by $\Delta_{s}$ with $s=1,\dots,S$. For each $\Delta_{s}$ we consider the piece-wise constant approximation of the intensity function $h$ introduced in Section 2
$h_{S}(x)=\frac{1}{\mu(\Delta_{s})}\int_{\Delta_{s}}h(t)\mu(dt),\quad x\in\Delta_{s},$ (6)
which is well-defined since $h$ is Lebesgue integrable on $K$. By Assumption 1, $h$ is continuous on $K$ a.e., thus, it is also uniformly continuous on $K$ a.e. As a consequence, let $\epsilon>0$, for $S$ large enough it holds
$\sup_{x\in K}$ $|h(x)-h_{S}(x)|=\max_{s=1,\dots,S}\sup_{x\in\Delta_{s}}|h(x)-h_{S}(x)|$
$\leq\max_{s=1,\dots,S}\Big{\{}\frac{1}{\mu(\Delta_{s})}\int_{\Delta_{s}}\sup_{x\in\Delta_{s}}|h(x)-h(t)|\mu(dt)\Big{\}}<\epsilon,\quad\mathrm{a.e.}$
Such piece-wise constant approximation holds on the whole $\mathbb{R}^{D}$ as it can be covered by an increasing sequence of compact sets $\{K_{m}\}_{m\geq 1}$ s.t. $\bigcup_{m=1}^{\infty}K_{m}=\mathbb{R}^{D}$. Therefore, $h_{S}\to h$ uniformly a.e. as $S\to\infty$. Now, we can choose such shrinking sub-regions as the closure of $B(x_{i},r_{i,k_{B,i}^{*}})$, because owing to Lemma 1 we can apply Lemma 2.3 of Biau and Devroye, (2015) to obtain that for all $i=1,\dots,n$, $t_{i,B}(k_{i}^{*})$ and $t_{i,A}(k_{i}^{*})$ tend to 0 a.s. as $n\to\infty$. To conclude we need to show that the distribution function of the data is proportional to the cumulative piece-wise constant intensity function of the Poisson process. Without loss of generality, we can set $S=n$ because as $n$ increases the number of balls increases and their size decreases. Given $h_{n}(x)\to h(x)$ uniformly a.e. as $n\to\infty$ and $|h_{n}(x)|\leq M$, owing to the Dominated Convergence Theorem it follows
[…]
Thus, letting $m\to\infty$ the cumulative intensity function based on $h_{n}$ converges to the true cumulative intensity function point-wise a.e. Consequently, let $F_{h}$ be the data distribution function based on $h$ and $F_{h_{n}}$ its approximation based on $h_{n}$, and note that again by the Dominated Convergence Theorem it follows
$\lim_{n\to\infty}F_{h_{n}}(t)$ $=\lim_{n\to\infty}\frac{1}{\int_{\mathbb{R}^{D}}h_{n}(x)\mu(dx)}\int_{-\infty}^{t}h_{n}(x)\mu(dx)$
$=\frac{1}{\int_{\mathbb{R}^{D}}h(x)\mu(dx)}\int_{-\infty}^{t}h(x)\mu(dx)=F_{h}(t),$
at every continuity point of $F_{h}$. Thus, we can invoke the Portmanteau theorem (Billingsley,, 1999) to obtain convergence in distribution, and the proof is concluded. ∎
Lemma 3.
For all $i=1,\dots,n$, the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$ are independent with increasing probability as $n\to\infty$.
Proof.
Consider the sets $B(x_{i},r_{i,k_{i,A}^{*}})$ and $B(x_{j},r_{j,k_{j,A}^{*}})$ for all $i\neq j$ and define the events
$E_{i,j}=\{B(x_{i},r_{i,k_{i,A}^{*}})\cap B(x_{j},r_{j,k_{j,A}^{*}})\neq\emptyset\}.$
We employ a counting argument, thus, let the total number of intersecting pairs be $W=\sum_{1\leq i<j\leq n}\mathbf{1}_{E_{i,j}}$ with expectation $\operatorname{E}[W]=\sum_{1\leq i<j\leq n}P(E_{i,j}).$ Now, note that for the volume of balls $B(x_{i},r_{i,k_{i,A}^{*}})$ it holds $\mu(B(x_{i},r_{i,k_{i,A}^{*}}))\asymp 1/n$ for all $i=1,\dots,n$ by the same arguments used in the proof of Lemma 1, and, thus, $r_{i,k_{i,A}^{*}}\asymp(1/n)^{1/d}$. Lemma 2 implies local uniformity on vanishing sets, therefore, the probability that two specific fixed balls intersect is proportional to the volume of the intersection region
$P(E_{i,j})$ $\asymp\mu\big{(}B(x_{i},r_{i,k_{i,A}^{*}})\cap B(x_{j},r_{j,k_{j,A}^{*}})\big{)}$
$\leq\min(\mu(B(x_{i},r_{i,k_{i,A}^{*}})),\mu(B(x_{j},r_{j,k_{j,A}^{*}})))\asymp 1/n$
and the expectation of $W$ is given by
$\operatorname{E}[W]=\sum_{1\leq i<j\leq n}P(E_{i,j})\lesssim\binom{n}{2}\frac{1}{n}=\frac{n-1}{2}.$
By a slight notation effort, we can extend this argument to $k$-wise intersections. For any $k=1,\dots,n$ we introduce indices $i_{1},\dots,i_{k}$ and and define the events
$E_{i_{1},\dots,i_{k}}=\Big{\{}\bigcap_{l=1}^{k}B(x_{i_{l}},r_{i_{l},k_{i_{l},A}^{*}})\neq\emptyset\Big{\}}.$
for all $\binom{n}{k}$ combinations without repetition of the indices $i_{1},\dots,i_{k}$. Clearly, when $k=1$ then $P(E_{i_{1}})=1$. Note that the probability of $k$-wise intersections can be bounded as follows
$P(E_{i_{1},\dots,i_{k}})$ $\asymp\mu\Big{(}\bigcap_{l=1}^{k}B(x_{i_{l}},r_{i_{l},k_{i_{l},A}^{*}})\Big{)}\leq\min_{l=1,\dots,k}\Big{\{}\mu(B(x_{i_{l}},r_{i_{l},k_{i_{l},A}^{*}}))\Big{\}}\asymp 1/n.$
Denote the number of $k$-wise intersections by $W_{k}=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\mathbf{1}_{E_{i_{1},\dots,i_{k}}}$. We can bound the expected value of $W_{k}$ as follows
$\operatorname{E}[W_{k}]=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}P(E_{i_{1},\dots,i_{k}})\lesssim\binom{n}{k}\frac{1}{n}\leq\frac{n^{k-1}}{k!}.$ (7)
Employing the notation introduced above, we can characterize the independence of the random variables $\{k_{A,j}^{*}|k_{B,j}^{*}\}_{j=1,\dots,n}$, for instance, through the event
$A_{n}=\bigcap_{k=1}^{n}\bigcap_{i_{1},\dots,i_{k}}{E}^{c}_{i_{1},\dots,i_{k}}$
and its limit $A:=\lim_{n\to\infty}A_{n}.$ In other words, we are enforcing that all the $k$-tuples of balls have empty intersections. It is immediate to note that for every sample size, the event $A$ has probability 1, however, we are interested in the following event
$A_{*}:=\{\nexists\>k\in\mathbb{N},i_{1},\dots,i_{k}:P(E_{i_{1},\dots,i_{k}})=1\},$
defined on the infinite-dimensional product probability space on which the collection of random vectors $X_{1},\dots,X_{\infty}$ is defined. Consider the empirical measure
$P_{n}(A_{*}^{c})=\frac{1}{\sum_{k=1}^{n}\binom{n}{k}}\sum_{k=1}^{n}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\mathbf{1}_{E_{i_{1},\dots,i_{k}}}=\frac{1}{2^{n}-1}\sum_{k=1}^{n}W_{k}$ (8)
and recall that, by the Law of Large Numbers, it converges to the true probability measure a.s. Now observe that, for a finite sample of size $n$, the index $k$ in principle ranges in $\{1,\dots,n\}$, however, we can show that its asymptotic behaviour can be linked to the set of random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$, allowing us to neglect a subset of indices for which intersections do not occur. To this aim, for any two balls centered at $x_{i}$ and $x_{j}$, and with radii $r_{i}$ and $r_{j}$, recall the intersection condition given by $\|x_{i}-x_{j}\|\leq r_{i}+r_{j}$. Hence, the set $\bigcap_{l=1}^{k}B(x_{i_{l}},r_{i_{l},k_{i_{l},A}^{*}})$ is non-empty whenever $i_{1},\dots,i_{k}$ are such that for all $i_{l}\subset\{i_{1},\dots,i_{k}\}$ it holds
$\|x_{i_{l}}-x_{i_{m}}\|\leq r_{i_{l},k_{i_{l},A}^{*}}+r_{i_{m},k_{i_{m},A}^{*}}$
for every other $i_{m}\subset\{i_{1},\dots,i_{k}\}$ with $i_{l}\neq i_{m}$. Thus, at least intuitively, such indices must be chosen such that points $x_{i_{1}},\dots,x_{i_{k}}$ are (uniformly) close enough according to the Euclidean metric. Resorting to the piece-wise approximation argument provided in Lemma 2, we consider an arbitrarily small compact subset $K^{*}_{\varepsilon}$ of the support with fixed measure $\mu(K^{*}_{\varepsilon})=\varepsilon$, such that the local intensity of the Poisson process achieves its maximum on average
$K^{*}_{\varepsilon}:=\arg\max_{K\in R^{D}:\mu(K)=\varepsilon}\frac{1}{\varepsilon}\int_{K}h(t)\mu(dt).$
Note that the largest number of intersecting balls have an intersection on $K^{*}_{\varepsilon}$ with high probability since, by definition, the number of points in $K^{*}_{\varepsilon}$, denoted by $N(K^{*}_{\varepsilon})$, is Poisson distributed with parameter
$\operatorname{E}[N(K^{*}_{\varepsilon})]=\int_{K^{*}_{\varepsilon}}h(t)\mu(dt)\geq\int_{K_{\varepsilon}}h(t)\mu(dt)=\operatorname{E}[N(K_{\varepsilon})]\quad\forall\,K_{\varepsilon}\in\mathbb{R}^{D},$
and the probability of $k$-wise intersections on $K^{*}_{\varepsilon}$ is proportional to $\operatorname{E}[N(K^{*}_{\varepsilon})]^{k}$. To bound $k$ it suffices to note that, for any $\varepsilon$ and $n$ large enough, there exists $X_{i}=x_{i}$ such that $x_{i}\in K^{*}_{\varepsilon}$ with associated $k_{A,i}^{*}$. Note that the largest $\varepsilon$ such that the intensity is constant on $K^{*}_{\varepsilon}$ makes the distance among points decrease uniformly, i.e. there exists $n$ large enough such that
$\sup_{x_{i},x_{j}\in K_{\varepsilon}^{*}}\|x_{i}-x_{j}\|\leq\inf_{i,j:x_{i},x_{j}\in K_{\varepsilon}^{*}}(r_{i,k_{A,i}^{*}}+r_{j,k_{A,j}^{*}}).$
Thus, to have the largest number of intersecting balls it must hold $\varepsilon\asymp 2r_{i,k_{A,i}^{*}}$ and, as a consequence, there exists $c_{1}\ll n$ such that intersections occur only for
$k\leq\max_{i=1,\dots n}\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}+c_{1}=:\widetilde{k}_{A}^{*}.$ (9)
To finally evaluate (8), we employ a martingale concentration argument to study the fluctuations of $W_{k}$ around $\operatorname{E}[W_{k}]$. Let $\{E_{j}\}_{j\in\mathbb{N}_{0}}$ be the sequence of events $E_{i_{1},\dots,i_{k}}$ sorted in lexicographical order with respect to the indices $i_{1},\dots,i_{k}$. We define the sequence $M_{t}=\sum_{j=0}^{t}\mathbf{1}_{E_{j}}-P(E_{j})$ and let $\mathcal{F}_{t}$ be the $\sigma$-field generated by the first $t$ random variables. Since $\operatorname{E}[M_{t+1}|\mathcal{F}_{t}]=M_{t}$, the sequence is a martingale with respect to the natural filtration. Moreover, it is immediate to notice that it possesses bounded increments, i.e. for any $t\in\mathbb{N}_{0}$ we have $|M_{t+1}-M_{t}|\leq 1$. Thus, by Azuma-Hoeffding inequality (Azuma,, 1967) for all $\epsilon>0$ and $t\in\mathbb{N}_{0}$
$P(|M_{t}-M_{0}|\geq\epsilon)\leq 2\exp\Big{(}-\frac{\epsilon^{2}}{2t}\Big{)}.$
Note that $M_{0}=0$ and the process $\{M_{t}\}_{t\in\mathbb{N}_{0}}$ tends to $W_{k}-\operatorname{E}[W_{k}]$, thus, since $t\leq\binom{n}{k}-1$ we obtain that, on an event with probability at least $1-\gamma$,
$|W_{k}-\operatorname{E}[W_{k}]|$ $<\sqrt{2t\log(2/\gamma)}\leq\sqrt{2\Big{(}\binom{n}{k}-1\Big{)}\log(2/\gamma)}$
$\leq\sqrt{\Big{(}\frac{2n^{k}}{k!}-2\Big{)}\log(2/\gamma)}\lesssim\frac{n^{k/2}}{\sqrt{k!}}\sqrt{\log(2/\gamma)}.$
As a consequence, for $\gamma=1/n$, the above inequality implies that the fluctuation of $W_{k}$ around its expectation is of order $O_{P}(\frac{n^{k/2}}{\sqrt{k!}}\sqrt{\log(2n)})$. Using this result together with (9), we obtain
$P_{n}(A_{*}^{c})$ $=\frac{1}{\sum_{k=1}^{n}\binom{n}{k}}\sum_{k=1}^{n}W_{k}=\frac{1}{2^{n}-1}\sum_{k=1}^{n}\Big{(}\operatorname{E}[W_{k}]+O_{P}\Big{(}\frac{n^{k/2}}{\sqrt{k!}}\sqrt{\log(2n)}\Big{)}\Big{)}$
$=\frac{1}{2^{n}-1}\sum_{k=1}^{\widetilde{k}_{A}^{*}}\Big{(}\operatorname{E}[W_{k}]+O_{P}\Big{(}\frac{n^{k/2}}{\sqrt{k!}}\sqrt{\log(2n)}\Big{)}\Big{)}$
$\sim\frac{1}{2^{n}-1}\Big{(}\operatorname{E}[W_{\widetilde{k}_{A}^{*}}]+O_{P}\Big{(}\frac{n^{\widetilde{k}_{A}^{*}/2}}{\sqrt{\widetilde{k}_{A}^{*}!}}\sqrt{\log(2n)}\Big{)}\Big{)}$
which converges to 0 as $n\to\infty$. Therefore, we have shown that the random vector $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$ is composed of $k_{A,i}^{*}$-dependent blocks which are negligible as $n\to\infty$. ∎
Lemma 4.
Let $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be the function given by $d\mapsto 0.2032^{1/d}$. The function $g$ is a contraction mapping on $(0,\infty)$, the function $\log\circ\,g$ is a contraction mapping on $(\sqrt{|\log(0.2032)|},\infty)$, and the function $\log(c_{2})/(\log\circ\,g)$ with $c_{2}\in\mathbb{R}$ is a contraction mapping on $(0,\infty)$ provided $c_{2}\in(0.2032,1/0.2032)$.
Proof.
Recall that a function $g$ is a contraction mapping on $(a,b)$ if its Lipschitz constant is smaller than 1 on $(a,b)$. Owing to the mean value theorem, it follows that the function $g$ is a contraction mapping on $(0,\infty)$ since
$\sup_{d\in(0,\infty)}|g^{\prime}(d)|=\sup_{d\in(0,\infty)}\Big{|}\frac{-g(d)\log(0.2032)}{d^{2}}\Big{|}\approx 1/3.$
The function $\log\circ\,g$ is not a contraction mapping on $(0,\infty)$, but since $|-d^{-2}{\log(0.2032)}|<1$ for $d>\sqrt{|\log(0.2032)|}\approx 1.26$, we have that $\log\circ\,g$ is a contraction mapping on $(\sqrt{|\log(0.2032)|},\infty)$. Finally, for $\log(c_{2})/(\log\circ\,g)$ we obtain
$\sup_{d\in(0,\infty)}$ $\Big{|}\frac{\partial}{\partial d}\frac{\log(c_{2})}{\log(g(d))}\Big{|}=\sup_{d\in(0,\infty)}\Big{|}\frac{\log(0.2032)\log(c_{2})}{d^{2}\log(g(d))^{2}}\Big{|}$
$=\sup_{d\in(0,\infty)}\Big{|}\frac{\log(0.2032)\log(c_{2})}{d^{2}(\frac{1}{d}\log(0.2032))^{2}}\Big{|}=\Big{|}\frac{\log(c_{2})}{\log(0.2032)}\Big{|},$
that is smaller than $1$ for $c_{2}\in(0.2032,1/0.2032)$, and the thesis follows. ∎
A.2 Algorithm convergence and asymptotic properties
We now prove the main results regarding the numerical convergence of Algorithm 1, as well as the consistency, and asymptotic normality of the proposed ID estimator. We start with the following proposition regarding the numerical convergence of Algorithm 1.
Proposition 1.
For a given tolerance $\delta>0$, there exists $n$ large enough such that, with high probability, Algorithm 1 converges.
Proof.
Let us consider the function $\tilde{g}:\mathbb{R}_{+}\to\mathbb{R}$ defined as
$d\mapsto\frac{\frac{1}{n}\sum_{i=1}^{n}{k_{A,i}^{*}(d)}}{\frac{1}{n}\sum_{i=1}^{n}{k_{B,i}^{*}(d)}}$
and note that Algorithm 1 is the fixed point iteration given by
$d_{m+1}=\frac{\log(\tilde{g}(d_{m}))}{\log(g(d_{m}))}:=G(d_{m})$ (10)
where $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is the function given by $d\mapsto 0.2032^{1/d}$ and $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$. The aim is to employ a (version of) the fixed point theorem to show that the sequence of iterations (10) converges to a fixed point. Let us consider a family of Borel sets $(\mathcal{G}_{h_{i}}^{i})_{h}\subset[0,\infty)$ given by
$\mathcal{G}_{h_{i}}^{i}=\{d\in[0,\infty):k_{i}^{*}(d)=h_{i}\},$
note that $(\mathcal{G}_{h_{i}}^{i})_{h_{i}}$ constitutes a partition of $[0,\infty)$. Now, let $d_{1}\neq d_{2}$, clearly in the trivial case where for all $i=1,\dots,n$ we have $d_{1},d_{2}\in\mathcal{G}_{h_{i}}^{i}$ then the map $G$ is a contraction mapping because it is constant. On the other hand, when $d_{1}\in\mathcal{G}_{h_{i}}^{i}$ and $d_{2}\in\mathcal{G}_{l_{i}}^{i}$ with $h_{i}\neq l_{i}$ we have that $G$ is not a contraction mapping in the usual sense because it is not continuous, therefore we require a different argument.
By Lemma 2, for every $i=1,\dots,n$ and as $n\to\infty$, the number of points in the hyper-sphere centered at $i$ of radius $r_{i,k_{B,i}^{*}}$ is Poisson distributed with parameter $\rho_{i}V_{i,k_{B,i}^{*}}$ while the number of points in the hyper-sphere centered at $i$ of radius $r_{i,k_{A,i}^{*}}$ is Poisson distributed with parameter $\rho_{i}V_{i,k_{A,i}^{*}}$. On the other hand, for every finite $n$, it is not restrictive to set $\rho_{A,i}V_{i,k_{A,i}^{*}}$ as the cumulative intensity on $B(x_{i},r_{i,k_{A,i}^{*}})$ with average intensity equal to $\rho_{i,A}$, and
$\rho_{B,i}V_{i,k_{B,i}^{*}}=\rho_{A,i}V_{i,k_{A,i}^{*}}+\rho_{B\setminus A,i}(V_{i,k_{B,i}^{*}}-V_{i,k_{A,i}^{*}})$
as the cumulative intensity on $B(x_{i},r_{i,k_{B,i}^{*}})$ with average intensity equal to $\rho_{B,i}$. Therefore, we have that
$\frac{\rho_{A,i}V_{i,k_{A,i}^{*}}}{\rho_{B,i}V_{i,k_{B,i}^{*}}}=\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}$ (11)
and $\rho_{A,i}/\rho_{B,i}=1$ whenever $\rho_{B\setminus A,i}=\rho_{A,i}$. By explicit computation of the conditional distribution, we have that
$p_{k_{A,i}^{*}|k_{B,i}^{*}}(x|y)$ $=\frac{p_{k_{A,i}^{*}}(x)p_{k_{B,i}^{*}-k_{A,i}^{*}}(y-x)}{p_{k_{B,i}^{*}}(y)}$
$=\frac{y!}{x!(y-x)!}\frac{(\rho_{A,i}V_{i,k_{A,i}^{*}})^{x}(\rho_{B\setminus A,i}(V_{i,k_{B,i}^{*}}-V_{i,k_{A,i}^{*}}))^{y-x}(\rho_{B,i}V_{i,k_{B,i}^{*}})^{-y}}{\exp(\rho_{A,i}V_{i,k_{A,i}^{*}}+\rho_{B\setminus A,i}(V_{i,k_{B,i}^{*}}-V_{i,k_{A,i}^{*}})-\rho_{B,i}V_{i,k_{B,i}^{*}})}$
$=\binom{y}{x}\frac{(\rho_{A,i}V_{i,k_{A,i}^{*}})^{x}(\rho_{B,i}V_{i,k_{B,i}^{*}}-\rho_{A,i}V_{i,k_{A,i}^{*}})^{y-x}}{(\rho_{B,i}V_{i,k_{B,i}^{*}})^{y}}\frac{(\rho_{B,i}V_{i,k_{B,i}^{*}})^{-x}}{(\rho_{B,i}V_{i,k_{B,i}^{*}})^{-x}}$
$=\binom{y}{x}\Big{(}\frac{\rho_{A,i}V_{i,k_{A,i}^{*}}}{\rho_{B,i}V_{i,k_{B,i}^{*}}}\Big{)}^{x}\Big{(}1-\frac{\rho_{A,i}V_{i,k_{A,i}^{*}}}{\rho_{B,i}V_{i,k_{B,i}^{*}}}\Big{)}^{y-x}$
Using (11) and conditioning on $\tau$ we get
$p_{k_{A,i}^{*}|k_{B,i}^{*},\tau}(x|y,\tau)=\binom{y}{x}\Big{(}\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}\Big{)}^{x}\Big{(}1-\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}\Big{)}^{y-x},$
that is
$k_{A,i}^{*}|k_{B,i}^{*},\tau(d)\sim\mathrm{Binomial}\Big{(}k^{*}_{B,i}(d),\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}\Big{)}.$
Without loss of generality, let us assume $d_{1}>d_{2}$, thus, it holds $k^{*}_{B,i}(d_{2})=k^{*}_{B,i}(d_{1})+c_{3}$ with $c_{3}\in\mathbb{N}$. Now, for every $i=1,\dots,n$, owing to Hoeffding’s inequality for bounded random variables (Boucheron et al.,, 2013) it follows
$P\Big{(}\Big{|}k_{A,i}^{*}|k_{B,i}^{*},\tau(d_{2})-k^{*}_{B,i}(d_{2})\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}\Big{|}\geq t\Big{)}\leq 2\exp\Big{(}-{\frac{2t^{2}}{{k^{*}_{B,i}(d_{2})}}}\Big{)}.$
By Lemma 3, the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$ are independent with increasing probability, that is they are independent up to an event of decreasing probability which is negligible for $n$ large enough, thus, we consider them as independent. As a consequence, averaging over $i$ and applying Hoeffding’s inequality again we obtain
$P\Big{(}\Big{|}\frac{1}{n}\sum_{i=1}^{n}k_{A,i}^{*}|k_{B,i}^{*},\tau(d_{2})-\frac{1}{n}\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}\Big{|}\geq t\Big{)}\leq 2\exp\Big{(}-{\frac{2n^{2}t^{2}}{\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})^{2}}}\Big{)}.$
Setting $\gamma=2\exp\Big{(}-{\frac{2n^{2}t^{2}}{\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})^{2}}}\Big{)}$ and solving for $t$ we obtain that for any $\gamma>0$, the following bound holds on an event with probability at least $1-\gamma$
$\Big{|}\frac{1}{n}\sum_{i=1}^{n}k_{A,i}^{*}|k_{B,i}^{*},\tau(d_{2})-\frac{1}{n}\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})\tau^{d}\frac{\rho_{i,A}}{\rho_{B,i}}\Big{|}<\sqrt{\frac{\log(2/\gamma)\sum_{i=1}^{n}k_{B,i}^{*}(d_{2})^{2}}{2n^{2}}}.$
By Lemma 1, we have that $\sum_{i=1}^{n}k_{B,i}^{*}(d_{2})^{2}\asymp n$, thus, as $n\to\infty$ we can choose $\gamma=1/n$, and obtain that $\frac{1}{n}\sum_{i=1}^{n}k_{A,i}^{*}(d_{2})$ belongs to a $O_{P}(\sqrt{\log(2n)/n})$-wide neighbourhood of its expectation. For $n$ large enough, it is not restrictive to assume ${\rho_{A,i}}/{\rho_{B,i}}$ is bounded above a finite constant and bounded away from 0, denoting the rounding function by $x\mapsto[x]$ and recalling that $\tau:=g(d_{2})$, as $n\to\infty$ we have
$\tilde{g}(d_{2})$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}k_{A,i}^{*}|k_{B,i}^{*},\tau(d_{2})}{\frac{1}{n}\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})}$
$=\frac{[\frac{1}{n}\sum_{i=1}^{n}k^{*}_{B,i}(d_{2})\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}+O_{P}(\sqrt{\log(2n)/n})]}{\frac{1}{n}\sum_{i=1}^{n}(k^{*}_{B,i}(d_{1})+c_{3})}$
$=\frac{[\frac{1}{n}\sum_{i=1}^{n}(k^{*}_{B,i}(d_{1})+c_{3})\tau^{d}\frac{\rho_{A,i}}{\rho_{B,i}}+O_{P}(\sqrt{\log(2n)/n})]}{\frac{1}{n}\sum_{i=1}^{n}(k^{*}_{B,i}(d_{1})+c_{3})}$
$=c_{2}\Big{(}g(d_{2})^{d}+\frac{O_{P}(\sqrt{\log(2n)/n})}{\frac{1}{n}\sum_{i=1}^{n}(k^{*}_{B,i}(d_{1})+c_{3})}\Big{)}\sim c_{2}g(d_{2})^{d}$
where $c_{2}$ is a positive constant. Therefore, $\log\circ\,\tilde{g}$ possesses the same asymptotic behaviour of $\log(c_{2})+d(\log\circ\,g)$, therefore, by Lemma 4, the function
$G=\frac{\log\circ\,\tilde{g}}{\log\circ\,{g}}=d+\frac{\log c_{2}}{\log\circ\,g}$
is a contraction mapping provided $c_{2}\in(0.2032,1/0.2032)$. Moreover, for every $\epsilon>0$ and for any $d_{2}$ it holds
$P\Big{(}\Big{|}\frac{\log(\tilde{g}(d_{2}))}{\log(g(d_{2}))}-d\Big{|}\leq\epsilon\Big{)}\to 1.$
Consequently, for every $d_{1}\neq d_{2}$ there exists $n$ large enough such that, with high probability, the inequality
$|G(d_{1})-G(d_{2})|=\Big{|}\frac{\log c_{2}}{\log g(d_{1})}-\frac{\log c_{2}}{\log g(d_{2})}\Big{|}\leq c_{4}|d_{1}-d_{2}|,\quad c_{4}\in[0,1)$
holds, because $c_{2}\to 1$ as a consequence of Lemma 2, and therefore ${\rho_{A,i}}/{\rho_{B,i}}\to 1$ for all $i=1,\dots,n$. Thus, as $n\to\infty$, $G$ converges uniformly to a constant contraction mapping with fixed point equal to $d$. In other words, for any given tolerance $\delta>0$, there exists $n$ large enough such that $\sup_{x\geq 0}|G(x)-d|<\delta$, that is $G$ is $\delta$-close to a contraction mapping. We obtain the desired result concerning the convergence of Algorithm 1 by employing (a version of) the fixed point theorem (see Boyd and Wong,, 1969), which implies that the sequence (10) converges with high probability at any given tolerance for $n$ large enough. Indeed, several versions of the fixed point theorem could be adopted, e.g. the fixed point theorem for quasi-contractions (Ćirić,, 1974), and others (see Kirk and Sims,, 2002). ∎
Remark 1.
Since $G$ converges to a constant function, we expect that 2 iterations are enough to obtain stabilization of Algorithm 1 at a given tolerance, and this is confirmed by the numerical experiments presented in the paper.
Note that the regularity of the intensity of the Poisson process can improve the convergence of the algorithm. In other words, for $\alpha$ not too small, whenever $h$ can be well approximated by a constant function on $B(x_{i},r_{i,k_{B,i}^{*}})$, i.e. when $\rho_{B\setminus A,i}\approx\rho_{A,i}$ for all $i=1,\dots,n$, convergence of Algorithm 1 at a given tolerance is guaranteed for much smaller sample sizes. In fact, in that case, we have an approximately uniform density on $B(x_{i},r_{i,k_{B,i}^{*}})$ implying that for every $d_{1},d_{2}\geq 0$ we have
$\tilde{g}(d_{2})=\frac{[{g}(d_{2})^{d}\frac{1}{n}\sum_{i=1}^{n}(k_{B,i}^{*}(d_{1})+c_{3})]}{\frac{1}{n}\sum_{i=1}^{n}(k_{B,i}^{*}(d_{1})+c_{3})}=c_{2}g(d_{2})^{d}$ (12)
where $c_{2}>0$ takes into account the approximation due to the rounding function and due to the local uniformity approximation. Hence, (12) implies that
$G(d_{2})=\frac{\log c_{2}+d\log(g(d_{2}))}{\log(g(d_{2}))}=\frac{\log c_{2}}{\log(g(d_{2}))}+d,$
and, thanks to Lemma 4, the function $d\mapsto{\log(c_{2})}/{\log(g(d))}$ is a contraction mapping provided $c_{2}\in(0.2032,1/0.2032)$, that is $c_{2}$ is not too far from $1$, and therefore, the rounding approximation error and the local uniformity approximation error are not too large. It is worth noticing that the arguments provided above prove that in general $\tilde{g}(x)=c_{2}g(x)^{d}$ where $c_{2}>0$ takes into account approximations errors mainly due to deviations from local uniformity. Moreover, the convergence of Algorithm 1 occurs when $c_{2}$ is not too large or too small. Now we derive consistency for the estimator $d^{*}_{n}$.
Proposition 2.
As $n\to\infty$, $d^{*}_{n}$ converges to $d$ in probability.
Proof.
As a consequence of the numerical convergence of Algorithm 1 derived in Proposition 1, it follows that for any $\epsilon>0$ and for any tolerance $\delta>0$, as $n\to\infty$
$P\Big{(}|d_{n}^{*}-d|>\epsilon\Big{)}$ $=P\Big{(}\bigcap_{\xi\in(-\delta,\delta)}\{|G(d_{n}^{*})+\xi-d|>\epsilon\}\Big{)}$
$\leq P\Big{(}\{|G(d_{n}^{*})+\delta-d|>\epsilon\}\Big{)}$
$=P\Big{(}\{|G(d_{n}^{*})-d|>\epsilon\}\Big{)}+o(1)$
$=P\Big{(}\Big{|}\frac{\log(\tilde{g}(d_{n}^{*}))}{\log(g(d_{n}^{*}))}-d\Big{|}>\epsilon\Big{)}+o(1)$
$=P\Big{(}\Big{|}\frac{\log(c_{2})+d\log({g}(d_{n}^{*}))}{\log(g(d_{n}^{*}))}-d\Big{|}>\epsilon\Big{)}+o(1)$
$=P\Big{(}\Big{|}\frac{\log(c_{2})}{\log(g(d_{n}^{*}))}\Big{|}>\epsilon\Big{)}+o(1).$
We conclude by noticing that $c_{2}\to 1$ by Lemma 2, thus, the upper bound tends to 0 as $n\to\infty$. ∎
The next theoretical result concerns the asymptotic normality of $d^{*}_{n}$. Indeed, the result follows by the properties of maximum likelihood estimators combined with numerical convergence of Algorithm 1 derived in Proposition 1.
Proposition 3.
There exists $c_{5}\in(0,1]$ such that the sequence $Z_{n}=\sqrt{nc_{5}I(d^{*}_{n})}(d^{*}_{n}-d)$ converges to $Z\sim\mathcal{N}(0,1)$ in distribution as $n\to\infty$.
Proof.
First observe that Algorithm 1 is an iterative conditional maximum likelihood estimator based on the conditional likelihood function
$L(d|k_{B,1}^{*},\dots,k_{B,n}^{*},\tau)=\prod_{i=1}^{n}\binom{k^{*}_{B,i}}{{k}_{A,i}^{*}}(\tau^{d})^{{k}_{A,i}^{*}}(1-\tau^{d})^{k^{*}_{B,i}-{k}_{A,i}^{*}}$
where ${k}_{A,i}^{*}=\sum_{j=1}^{n-1}\mathbf{1}\{\tau r_{ik_{B,i}^{*}}-r_{i,j}>0\}$. Moreover, by Lemma 2 and Assumption 1 it follows that for every $\alpha>0$, $P(k^{*}_{i}:\rho_{i}=\rho_{i}^{\prime})\to 1$ for every $i=1,\dots,n$. Thanks to Proposition 1 we know that Algorithm 1 is a fixed point iteration that converges to its fixed point. As $n\to\infty$ such fixed point is equal to $d$ since ${\rho_{A,i}}/{\rho_{B,i}}\to 1$ for every $i=1,\dots,n$ as a consequence of Assumption 1. Therefore, $L(d|k_{B,1}^{*},\dots,k_{B,n}^{*},\tau)$ is (asymptotically) an ordinary likelihood function because $k_{B,1}^{*},\dots,k_{B,n}^{*}$ and $\tau$ converge with probability tending to 1. Furthermore, the estimation problem is asymptotically regular since the distribution $\mathrm{Binomial}(k_{B,i}^{*},\tau^{d})$ with fixed $k_{B,i}^{*}$ belongs to the exponential family. In conclusion, the asymptotic normality of $Z_{n}$ follows as a direct consequence of standard asymptotic properties of M-estimators Van der Vaart, (2000). In particular, the thesis follows employing the usual asymptotic linearity property of the maximum likelihood estimator (Theorem 5.39 of Van der Vaart,, 2000) together with a version of the Central Limit Theorem for $m$-dependent random variables (see e.g. Berk,, 1973, Lehmann,, 1999, Janson,, 2021), whose conditions are satisfied as a consequence of boundedness of the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$, and since
$\frac{1}{n}\operatorname{Var}$ $\Big{[}\sum_{i=1}^{n}k_{A,i}^{*}|k_{B,i}^{*},\tau\Big{]}=\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}[k_{A,i}^{*}|k_{B,i}^{*},\tau]+\frac{1}{n}\sum_{i=1}^{n}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\operatorname{Cov}[k_{A,i}^{*}|k_{B,i}^{*},\tau;k_{A,j}^{*}|k_{B,j}^{*},\tau]$
$\lesssim\frac{1}{4n}\sum_{i=1}^{n}k_{B,i}^{*}+2\max_{i=1,\dots,n}\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}\max_{\begin{subarray}{c}i,j=1,\dots,n\\ j\neq i\end{subarray}}\{|\operatorname{Cov}[k_{A,i}^{*}|k_{B,i}^{*},\tau;k_{A,j}^{*}|k_{B,j}^{*},\tau]|\}$
$\leq\frac{1}{4n}\sum_{i=1}^{n}k_{B,i}^{*}+2\max_{i=1,\dots,n}\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}\max_{\begin{subarray}{c}i,j=1,\dots,n\\ j\neq i\end{subarray}}\{\operatorname{Var}[k_{A,i}^{*}|k_{B,i}^{*},\tau]^{\frac{1}{2}}\operatorname{Var}[k_{A,j}^{*}|k_{B,j}^{*},\tau]^{\frac{1}{2}}\}$
$\leq\frac{1}{4n}\sum_{i=1}^{n}k_{B,i}^{*}+2\max_{i=1,\dots,n}\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}\max_{\begin{subarray}{c}i,j=1,\dots,n\\ j\neq i\end{subarray}}\{\frac{k_{B,i}^{*}k_{B,j}^{*}}{4}\}<\infty.$
Notice that, in the previous chain of inequalities, we used the binomiality of the random variables $\{k_{A,i}^{*}|k_{B,i}^{*},\tau\}_{i=1,\dots,n}$, Cauchy-Schwartz inequality and Lemma 1. The proof is concluded. ∎
Remark 2.
Numerical experiments show that the variance inflating constant $c_{5}$ (due to $k_{A,i}^{*}$-dependent blocks) can be set equal to 1 with no significant loss of accuracy when approximating the limiting distribution, see Figure 8.
Appendix B Bayesian version of ABIDE
The Bayesian version of the proposed approach is a straightforward extension of the Bayesian estimator presented in the supplementary material of Macocco et al., (2023). Here we briefly outline it and show how the adaptive version can be obtained. By a conjugacy argument, consider the following prior, $p=\tau^{d}\sim\mathrm{Beta}(\alpha_{0},\beta_{0})$, and derive the corresponding posterior distribution
$p|k_{B,1},\dots,k_{B,n}\sim\mathrm{Beta}(\alpha,\beta)$
where $\alpha=\alpha_{0}+\sum_{i=1}^{n}k_{A,i}$ and $\beta=\beta_{0}+\sum_{i=1}^{n}(k_{B,i}-k_{A,i})$. In our adaptive framework, we have instead the following posterior distribution
$p^{*}|k_{B,1},\dots,k_{B,n}\sim\mathrm{Beta}(\alpha^{*},\beta^{*})$
where $\alpha^{*}=\alpha_{0}+\sum_{i=1}^{n}k_{A,i}^{*}$ and $\beta^{*}=\beta_{0}+\sum_{i=1}^{n}(k_{B,i}^{*}-k_{A,i}^{*})$. Therefore, omitting the conditioning on the $k_{B,1},\dots,k_{B,n}$ for notational convenience, the posterior density of the ID, derived by a change of variable argument, is given by
$f_{d}(x)=|\tau^{x}\log(\tau)|\tau^{x(\alpha^{*}-1)}(1-\tau^{x})^{\beta^{*}-1}$
where the posterior expectation and the posterior variance can be derived using the well-known fact that for $X\sim\mathrm{Beta}(\alpha,\beta)$ we have
$\operatorname{E}[X]=\psi_{0}(\alpha)-\psi_{0}(\alpha+\beta),\quad\operatorname{Var}[X]={\psi_{1}(\alpha)-\psi_{1}(\alpha+\beta)}$
where $\psi_{0}(z)=\frac{d}{dz}\log\Gamma(z)$ and $\psi_{1}(z)=\frac{d^{2}}{dz^{2}}\log\Gamma(z)$ are the digamma and the trigamma functions respectively. Thus, we obtain the posterior expectation and the posterior variance
$\operatorname{E}[d]=\frac{\psi_{0}(\alpha^{*})-\psi_{0}(\alpha^{*}+\beta^{*})}{\log(\tau)},\quad\operatorname{Var}[d]=\frac{\psi_{1}(\alpha^{*})-\psi_{1}(\alpha^{*}+\beta^{*})}{\log(\tau)^{2}}.$ (13)
Therefore, adopting a squared error loss function, $\operatorname{E}[d]$ is a proper Bayesian estimator of the ID and, it gives rise to a Bayesian version of Algorithm 1 obtained setting $d_{\text{next}}=\operatorname{E}[d]$. This is named the Bayesian Adaptive Binomial ID Estimator (BABIDE).
Proposition 4.
BABIDE is asymptotically equivalent to ABIDE and inherits the same theoretical guarantees.
Proof.
Using the well-known asymptotic approximation of $\psi_{0}$ given by
$\psi_{0}(z)\sim\log(z)-\sum_{j=1}^{\infty}\frac{\zeta(1-j)}{z^{j}}$
where $\zeta$ is the Riemann Zeta function, we obtain the asymptotic equivalence
$\operatorname{E}[d]\sim\frac{\log(\sum_{i=1}^{n}k_{A,i}^{*})-\log(\sum_{i=1}^{n}k_{B,i}^{*})}{\log(\tau)},$
and, as a consequence, numerical convergence of BABIDE is naturally inherited by Proposition 1. Moreover, note that statistical guarantees regarding consistency and asymptotic normality derived for ABIDE can be directly extended for the Bayesian version by an immediate posterior consistency argument relying on the analytical expression of the posterior variance in (13), and using the classical Bernstein–von Mises theorem (Le Cam,, 1986) for the posterior distribution. ∎
Appendix C Adaptive Generalized Ratios ID Estimator
Another approach that fits into the same theoretical framework outlined in the paper is proposed in Denti et al., (2022). Let $r_{i,l}$ be the distance between unit $i$ and its $l$-th NN according to a selected metric. Then, let $n_{1},n_{2}$ be two positive integers with $n_{2}>n_{1}$ and define the ratio $\mu_{i,n_{1},n_{2}}=r_{i,n_{2}}/r_{i,n_{1}}$ which is a.s. well-defined for continuous data. For discrete data, duplicated points leading to zero values at the denominator of $\mu_{i}$, need to be removed. The authors derive the density function of $\mu_{i,n_{1},n_{2}}$:
$f_{\mu_{i,n_{1},n_{2}}}(\mu)=\frac{d(\mu^{d}-1)^{n_{2}-n_{1}-1}}{\mu^{d(n_{2}-1)+1}\mathrm{B}(n_{2}-n_{1},n_{1})},\quad\mu>1$ (14)
where $\mathrm{B}(\cdot,\cdot)$ is the Beta function and $d$ is the ID. In the following, we omit $n_{1}$ and $n_{2}$ from the subscript for notational convenience. Assuming independence of $\mu_{1},\dots,\mu_{n}$, the Authors introduce the Generalized Ratios ID Estimator (GRIDE) given by
$\begin{split}\widehat{d}&=\arg\underset{d\geq 0}{\max}\,\log L(d|\mu_{1},\dots,\mu_{n})\\ &=\arg\underset{d\geq 0}{\max}\,\log\prod_{i=1}^{n}\frac{d(\mu_{i}^{d}-1)^{n_{2}-n_{1}-1}}{\mu_{i}^{d(n_{2}-1)+1}\mathrm{B}(n_{2}-n_{1},n_{1})},\end{split}$ (15)
where the maximization is carried out numerically. This approach generalizes the 2NN estimator proposed in Facco et al., (2017), since when $n_{1}=1$ and $n_{2}=2$, the density in (14) reduces to a $\mathrm{Pareto}(1,d)$ and the unbiased maximum likelihood estimator is
$\widehat{d}=\frac{n-1}{\sum_{i=1}^{n}\log(\mu_{i})},$
equivalent to the estimator proposed in Facco et al., (2017). It must be pointed out that, when compared to the GRIDE estimator based on values of $n_{2}>2$, the 2NN is more likely to satisfy the local homogeneity assumption of the Poisson process density because it only uses information up to the scale of the second NN of each statistical unit. On the other hand, it is potentially affected by the presence of noise in the data. The selection of $n_{1}$ and $n_{2}$ is somehow linked to the selection of the radii in the framework of Macocco et al., (2023), where the choice of the radii induce $k_{A,1}\dots,k_{A,n},k_{B,1}\dots,k_{B,n}$ in estimator (1), and this choice is affected by the same trade-off between robustness to noise and local homogeneity. The estimator in (15) is straightforwardly extended to its adaptive version by splitting the (log)likelihood in its single unit contributions, plugging in the unit dependent $n_{2,i}=k^{*}_{i}$ and, as suggested in Denti et al., (2022), taking $n_{1,i}=n_{2,i}/2$. This leads to an iterative algorithm similar to ABIDE. We might refer to it as Adaptive GRIDE (AGRIDE). Some simulations showed a comparable behaviour of ABIDE and AGRIDE. However, it should be recalled that ABIDE is based on BIDE, which has a closed-form expression and does not require additional numerical routines to be computed. This is a relevant advantage as it makes the adaptive version mathematically more convenient when deriving theoretical guarantees.
Appendix D Additional experiments
Figure 6 reports a comparison of the distributions of $k^{*}_{i}$ (left) and corresponding radii $t_{B,i}(k_{i}^{*})$ (right) obtained using the ID estimated with the 2NN estimator (blue) against those obtained with ABIDE (orange) in reference to Panel E of Figure 3. The associated IDs are $\approx 5$ for 2NN and $\approx 2.35$ for ABIDE. One can appreciate how the neighbourhoods become larger together with the lowering of the ID. This shows how a proper estimation of the ID also affects, non-trivially, the estimation of $k^{*}_{i}$ and vice-versa. The right panel displays the distribution of the distances associated with the selected neighbours. One can appreciate how also this distribution is broad, with a long tail and two distinct peaks at small distances. The shape of such distributions allows us to understand why the results obtained at fixed uniform radius $t_{B,i}=t_{B}$ or at fixed uniform neighbourhood extension $k_{i}=k$ could not provide reliable results for this kind of dataset, where the density is rapidly changing in a non-uniform way across the dataset.
Figure 6 shows the evolution of $\overline{k^{*}}=\frac{1}{n}\sum_{i=1}^{n}k_{i}^{*}$ (top) and corresponding $\overline{t_{A}(k^{*})}=\frac{1}{n}\sum_{i=1}^{n}t_{A,i}(k_{i}^{*})$ (bottom) computed on multivariate Gaussian data. In each panel, the parameters that do not vary are fixed at $n=10,000$, $d=5$ and $\alpha=0.01$. It is at once apparent that $\overline{k^{*}}$ scales sub-linearly with the sample size tending to get concetrated on a finite compact set (top-left), and $\overline{t_{A}(k^{*})}$ decreases as $n$ increases (bottom-left), making the neighbourhoods asymptotically smaller and smaller, therefore, allowing us to consider the random variables $\{k_{A,i}|k_{B,i},\tau\}_{i=1,\dots,n}$ nearly independent (Lemma 3). In the central panels, we generate data with different ID values and compute $\overline{k^{*}}$ and $\overline{t_{A}(k^{*})}$ using the true ID. It is clear that $\overline{k^{*}}$ decreases as the ID increases since density variations happen at smaller neighbourhoods, while $\overline{t_{A}(k^{*})}$ increases because higher dimensions imply higher distance between data points. Finally, on the right panels, we show that a larger $\alpha$ leads to smaller neighbourhoods ensuring higher protection against type 2 errors.
Figure 8 reports the Monte Carlo distribution of the centred, re-scaled and normalized statistic $Z_{n}=\sqrt{nI(d^{*}_{n})}(d^{*}_{n}-d)$ for $n=10,000$ based on 500 Monte Carlo independent replicas of a 5-dimensional uniform distribution with periodic boundary conditions. From the figure it is at once apparent that the distribution of $Z_{n}$ tends to a standard normal, thus confirming the results of Proposition 3. Moreover, this experiment confirms that the effect of dependencies among $\{k_{A,i}|k_{B,i},\tau\}_{i=1,\dots,n}$ vanishes for large sample size, as also shown in Lemma 3.
Figure 6: Möbius dataset: comparison of the distributions of $k^{*}_{i}$ (left) and corresponding radii $t_{B,i}(k_{i}^{*})$ (right) obtained using the ID estimated with the 2NN estimator (blue) against those obtained with ABIDE (orange) in reference to Panel E of Figure 3.
Figure 7: Evolution of $\overline{k^{*}}=\frac{1}{n}\sum_{i=1}^{n}k_{i}^{*}$ (top) and corresponding $\overline{t_{A}(k^{*})}=\frac{1}{n}\sum_{i=1}^{n}t_{A,i}(k_{i}^{*})$ (bottom) computed on multivariate Gaussian data. In each panel, the parameters that do not vary are fixed at $n=10,000$, $d=5$ and $\alpha=0.01$.
Figure 8: Monte Carlo distribution of the centred, re-scaled and normalized statistic $Z_{n}=\sqrt{nI(d^{*}_{n})}(d^{*}_{n}-d)$ for $n=10,000$ based on 500 Monte Carlo independent replicas of a 5-dimensional uniform distribution with periodic boundary conditions.
Acknowledgements
A.M. acknowledges financial support from SNSF grant 208249. The views and opinions expressed in this paper are those of the authors and do not necessarily reflect the official policy or position of Banca d’Italia.